### 05 Mar 2026 v1

# Biquadratic SOS Rank and Double Zarankiewicz Number

Abstract
The maximum sum-of-squares (SOS) rank of an $m\times n$ biquadratic form, denoted $\operatorname{BSR}(m,n)$, has recently been shown to be bounded below by the classical Zarankiewicz number $z(m,n)$, the maximum number of edges in a $C_{4}$-free bipartite graph. However, a gap was discovered for the $4\times 3$ case: $\operatorname{BSR}(4,3)\geq 8>z(4,3)=7$, arising from a construction that adds a square of a bilinear form $(x_{4}y_{2}+x_{1}y_{3})^{2}$ to a $C_{4}$-free simple biquadratic form. This paper introduces the double Zarankiewicz number $z_{2}(m,n)$ to capture such phenomena. We define a new class of bipartite graphs that may contain both ordinary edges (1-edges) and “double edges” (2-edges), the latter corresponding to squares of simple two-term bilinear forms. A graph $G$ of this type yields a doubly simple biquadratic form $P_{G}$, and we prove that if $G$ contains no generalised $C_{4}$-cycle then $\operatorname{sos}(P_{G})=|E_{1}|+|E_{2}|$; consequently $\operatorname{BSR}(m,n)\geq z_{2}(m,n)$. We determine the exact value of $z_{2}(m,n)$ for several small parameters: $z_{2}(m,2)=m+1$, $z_{2}(2,n)=n+1$, $z_{2}(3,3)=6$, $z_{2}(4,3)=8$, and $z_{2}(5,3)=9$. For the $4\times 4$ case we establish $10\leq z_{2}(4,4)\leq 11$, leaving the exact value as an open problem. These results improve the known lower bounds for $\operatorname{BSR}(4,4)$ and $\operatorname{BSR}(5,3)$. The paper concludes with a discussion of open questions and connections between extremal graph theory and the study of SOS representations.
Keywords. Biquadratic form, sum of squares, SOS rank, Zarankiewicz number, double Zarankiewicz number, bipartite graph, $C_{4}$-cycle, extremal graph theory.
AMS subject classifications. 14P10, 05C35, 11E25, 15A69, 90C22.
1 Introduction
Denote $\{1,\ldots,m\}$ as $[m]$. Assume that $m,n\geq 2$. Let
$P(\mathbf{x},\mathbf{y})=\sum_{i,k=1}^{m}\sum_{j,l=1}^{n}a_{ijkl}x_{i}x_{k}y_{j}y_{l}.$ (1)
We call $P$ an $m\times n$ biquadratic form. Here $a_{ijkl}$ are real numbers. We assume that
$a_{ijkl}=a_{kjil}=a_{klij}$ (2)
for $i,k\in[m],j,l\in[n]$. A PSD (positive semi-definite) biquadratic form is one for which $P(\mathbf{x},\mathbf{y})\geq 0$ for all $\mathbf{x},\mathbf{y}$. It is SOS (sum of squares) if it can be written as a finite sum of squares of bilinear forms, i.e.,
$P({\bf x},{\bf y})=\sum_{p=1}^{r}f_{p}({\bf x},{\bf y})^{2},$ (3)
where $f_{p}$ are $m\times n$ bilinear forms for $p\in[r]$. The smallest $r$ such that (3) holds is called the SOS rank of $P$. In this case, we say that $\sum_{p=1}^{r}f_{p}({\bf x},{\bf y})^{2}$ is irreducible. The maximum SOS rank of $m\times n$ SOS biquadratic forms is denoted as $BSR(m,n)$. It is known that $BSR(m,2)=m+1$ and $BSR(2,n)=n+1$ [1], $BSR(3,3)=6$ [2] and $BSR(m,n)\leq mn-1$ [7].
Recently, it was proved [5] that
$BSR(m,n)\geq z(m,n),$ (4)
where $z(m,n)$ is the Zarankiewicz number [10]. The Zarankiewicz problem, originating from [10], asks for the maximum number of edges in an $m\times n$ bipartite graph that contains no complete bipartite subgraph $K_{2,2}$ (i.e., no $C_{4}$). This classical extremal problem has been extensively studied; see [6, 3, 8] for surveys and fundamental bounds.
The inequality (4) is based upon the simple biquadratic forms introduced in [7]. A biquadratic form is a simple biquadratic form if it contains only distinct terms of the type $x_{i}^{2}y_{j}^{2}$. Then we may establish a one-to-one relation. Let $G=(S,T,E)$ be a bipartite graph, where $S=\{1,\cdots,m\}$ and $T=\{1,\cdots,n\}$ are the vertex sets and $E$ is the edge set. Then uniquely we have a simple quadratic form
$P_{G}({\bf x},{\bf y})=\sum_{(i,j)\in E}x_{i}^{2}y_{j}^{2}.$ (5)
If $G$ has no $C_{4}$-cycles, then the simple biquadratic form $P_{G}$ is irreducible, i.e., $\mathrm{sos}(P)=|E|$. On the other hand, the Zarankiewicz number $z(m,n)$ is defined as the largest number of edges such that the $m\times n$ bipartite graph $G$ is $C_{4}$-cycle free [10, 8]. Thus, we have (4).
It is known that $BSR(m,2)=z(m,2)=m+1$, $BSR(2,n)=z(2,n)=n+1$ and $BSR(3,3)=z(3,3)=6$. However, it was found in [9] that $BSR(4,3)\geq 8>z(4,3)=7$. The value $z(4,3)=7$ is a classical result in Zarankiewicz theory [8, 6]. This gap was discovered in the following way. First, we have
$P_{4,3,7}({\bf x},{\bf y})=x_{1}^{2}y_{1}^{2}+x_{2}^{2}y_{2}^{2}+x_{3}^{2}y_{3}^{2}+x_{1}^{2}y_{2}^{2}+x_{2}^{2}y_{3}^{2}+x_{3}^{2}y_{1}^{2}+x_{4}^{2}y_{1}^{2}.$
This is a $4\times 3$ simple biquadratic form whose corresponding $4\times 3$ bipartite graph $G$ attains the Zarankiewicz number $z(4,3)=7.$ Then, an $8$-square $4\times 3$ sos biquadratic form $Q$ is formed as
$Q({\bf x},{\bf y})=P_{4,3,7}({\bf x},{\bf y})+(x_{4}y_{2}+x_{1}y_{3})^{2}.$
It was proved that $Q$ is irreducible, i.e., $\mathrm{sos}(Q)=8$. In this way, it was established that
$BSR(4,3)\geq 8>z(4,3)=7.$ (6)
This makes us think further: Why can such a thing happen? What is its meaning in the theory of the Zarankiewicz number? Now, $(x_{4}y_{2}+x_{1}y_{3})^{2}$ is not a single term square. It does not correspond to an edge of the bipartite graph $G$. But can we interpret it as a super edge of $G$? This motivates this paper.
The remainder of this paper is organized as follows. In Section 2, we introduce the concept of a double Zarankiewicz number $z_{2}(m,n)$ and establish its basic properties, including its relation to the classical Zarankiewicz number and exact values for small parameters. We also prove the fundamental inequality $\operatorname{BSR}(m,n)\geq z_{2}(m,n)$ by showing that any graph attaining the double Zarankiewicz number yields an irreducible doubly simple biquadratic form. Section 3 investigates the $4\times 4$ case, presenting a construction achieving ten edges and establishing the bounds $10\leq z_{2}(4,4)\leq 11$. Section 4 treats the $5\times 3$ case, where we prove $z_{2}(5,3)=9$. Finally, Section 5 concludes the paper with a summary of results and a discussion of open problems.
2 Double Zarankiewicz Number
Let $G=(S,T,E)$ be an $m\times n$ bipartite graph, where $S=\{1,\cdots,m\}$ and $T=\{1,\cdots,n\}$ are its vertex sets. We define the edge set $E=E_{1}\cup E_{2}$, where $E_{1}$ is the $1$-edge set, and $E_{2}$ is the $2$-edge set. A $1$-edge $e$ in $E_{1}$ has the form $e=(i,j)$, where $i\in S$ and $j\in T$. On the other hand, a $2$-edge $e$ in $E_{2}$ has the form $(i,j;k,l)$, where $i,k\in S$ and $j,l\in T$. It is fine that $i=k$ or $j=l$ but should not be both $i=k$ and $j=l$. For such a bipartite graph $G$, we associate it with an $m\times n$ sos biquadratic form $P_{G}$, defined as:
$P_{G}({\bf x},{\bf y})=\sum_{(i,j)\in E_{1}}x_{i}^{2}y_{j}^{2}+\sum_{(i,j;k,l)\in E_{2}}(x_{i}y_{j}+x_{k}y_{l})^{2}.$ (7)
We call such a biquadratic form a doubly simple biquadratic form. If $P_{G}$ is irreducible, then the sos rank of $P_{G}$ is $|E|=|E_{1}|+|E_{2}|$.
Let $m=4$ and $n=3$, $E_{1}=\{(1,1),(2,2),(3,3),(1,2),(2,3),(3,1),(4,1)$ and $E_{2}=\{(i,j;k,l)\}$. Then we see if $(i,j;k,l)$ is one of the following:
(4,2;1,3),(4,3;1,2),(4,2;4,3),(4,1;4,2),(4,2;1,3),\;(4,3;1,2),\;(4,2;4,3),\;(4,1;4,2),
$P_{G}$ is irreducible. We wish to characterize the combinatoric feature of $(i,j;k,l)$.
We now refine the notion of a $C_{4}$-cycle to account for the presence of 2-edges and formulate a linear independence condition that characterizes irreducibility.
Definition 2.1 (Double Zarankiewicz Number).
Let $G=(S,T,E_{1}\cup E_{2})$ be an $m\times n$ bipartite graph with vertex sets $S=[m]$, $T=[n]$, where
  • $E_{1}$ is a set of 1-edges of the form $(i,j)$ with $i\in S$, $j\in T$;
  • $E_{2}$ is a set of 2-edges of the form $(i,j;k,l)$ with $i,k\in S$ and $j,l\in T$ (allowing $i=k$ or $j=l$, but not both simultaneously).
Assume that there are no repeated 1-edges or 2-edges. We say that $G$ contains a generalized $C_{4}$-cycle if either
  1. (i)
    there exist distinct rows $i,k\in S$ and distinct columns $j,l\in T$ such that, counting each 1-edge as one ordinary edge and each 2-edge $(p,q;r,s)$ as two ordinary edges $(p,q)$ and $(r,s)$, the total number of ordinary edges (with multiplicities) contained in the set $\{i,k\}\times\{j,l\}$ is at least four (if the vertices of a 2-edge are not both in the block $\{i,k\}\times\{j,l\}$, then this 2-edge is not counted);
  2. (ii)
    there exists a nontrivial linear combination
    $\sum_{(p,q)\in E_{1}}\alpha_{pq}\,\mathbf{e}_{pq}\;+\!\!\sum_{(i,j;k,l)\in E_{2}}\beta_{ijkl}\,(\mathbf{e}_{ij}+\mathbf{e}_{kl})=\mathbf{0}$
    in $\mathbb{R}^{mn}$ (where $\mathbf{e}_{pq}$ is the standard basis vector corresponding to $x_{p}y_{q}$), with coefficients not all zero, and such that there exist distinct rows $r,s\in[m]$ and distinct columns $t,u\in[n]$ for which every basis vector $\mathbf{e}_{pq}$ appearing in the combination (i.e., with $\alpha_{pq}\neq 0$ or as part of a 2-edge term with $\beta_{ijkl}\neq 0$) satisfies $p\in\{r,s\}$ and $q\in\{t,u\}$.
The double Zarankiewicz number $z_{2}(m,n)$ is the maximum possible total number of edges $|E_{1}|+|E_{2}|$ for which such a generalized $C_{4}$-cycle does not exist.
Theorem 2.2.
The double Zarankiewicz number satisfies
$z_{2}(m,n)\geq z(m,n)\ \text{for all }m,n\geq 2,$
$z_{2}(m,2)=m+1\quad\text{for all }m\geq 2,$
$z_{2}(2,n)=n+1\quad\text{for all }n\geq 2,$
$z_{2}(3,3)=6,$
$z_{2}(4,3)\geq 8>z(4,3).$
Proof.
By definition, taking $|E_{2}|=0$ we have $z_{2}(m,n)\geq z(m,n)$ for all $m,n\geq 2$.
Now consider $(m,2)$. Let the two columns be $1$ and $2$. For any admissible graph $G=(S,T,E_{1}\cup E_{2})$ with $|T|=2$, define for each row $i\in S$ the total multiplicity $s_{i}$ of ordinary edges incident to that row, i.e., the number of times the cell $(i,1)$ appears plus the number for $(i,2)$, counting each 1-edge as one and each 2-edge $(p,q;r,s)$ as one ordinary edge for each of its two constituent cells. Then for any two distinct rows $i,k$, the submatrix $\{i,k\}\times\{1,2\}$ contains exactly $s_{i}+s_{k}$ ordinary edges. Condition (i) of Definition 2.1 requires $s_{i}+s_{k}\leq 3$ for every pair $i\neq k$; otherwise a $C_{4}$ would appear with at least four edges.
We now bound the total sum $M=\sum_{i=1}^{m}s_{i}$. If the largest $s_{i}$ equals $2$, then all other rows must have $s_{i}\leq 1$ (since $2+2=4$ is forbidden). Hence $M\leq 2+(m-1)\cdot 1=m+1$. If the largest equals $3$, then every other row must have $s_{i}=0$ (because $3+1=4$ would violate the condition), so $M\leq 3$. If all $s_{i}\leq 1$, then $M\leq m$. In all cases $M\leq m+1$.
The total number of edges $T=|E_{1}|+|E_{2}|$ satisfies $M=|E_{1}|+2|E_{2}|$ because each 2-edge contributes two ordinary edges. Hence $T=M-|E_{2}|\leq M\leq m+1$. Thus $z_{2}(m,2)\leq m+1$.
Conversely, the star graph where one row has both columns (i.e., two 1-edges) and every other row has a single 1-edge to column $1$ (or symmetrically) contains exactly $m+1$ edges and no $C_{4}$ (any two rows share at most one column). Hence it is admissible, giving $z_{2}(m,2)\geq m+1$. Therefore $z_{2}(m,2)=m+1$. By symmetry, interchanging rows and columns yields $z_{2}(2,n)=n+1$.
Now consider $(3,3)$. For a graph $G$ on $3$ rows and $3$ columns, there are $\binom{3}{2}^{2}=9$ choices of two distinct rows and two distinct columns. For each such $2\times 2$ submatrix, condition (i) limits the total number of ordinary edges (with multiplicities) contained in it to at most $3$. Summing over all nine submatrices we obtain a total $S\leq 9\times 3=27$.
Every ordinary edge $(i,j)$ (counted with its multiplicity) appears in exactly (31)(31)=4(3-1)(3-1)=4 of these submatrices (choose one other row and one other column). If $M$ denotes the total number of ordinary edges (i.e., $M=|E_{1}|+2|E_{2}|$), then $4M=S\leq 27$, so $M\leq 6.75$ and consequently $M\leq 6$. Since $T=|E_{1}|+|E_{2}|\leq M$, we obtain $T\leq 6$. Thus any admissible graph has at most $6$ edges. The Zarankiewicz number $z(3,3)=6$ is attained by a $C_{4}$-free bipartite graph (e.g., a $3\times 3$ grid with six edges); taking $E_{2}=\emptyset$ gives an admissible configuration with $6$ edges. Hence $z_{2}(3,3)=6$.
Finally, for $(4,3)$, we exhibit a construction with total edges $8$. Let
$E_{1}=\{(1,1),(2,2),(3,3),(1,2),(2,3),(3,1),(4,1)\},\hskip 28.80008ptE_{2}=\{(4,2;1,3)\}.$
We verify that this graph contains no generalized $C_{4}$-cycle according to Definition 2.1.
Condition (i). For each pair of distinct rows $i,k\in\{1,2,3,4\}$ and distinct columns $j,l\in\{1,2,3\}$, we count the number of 1?edges in $\{i,k\}\times\{j,l\}$ plus twice the number of 2-edges whose two constituent ordinary edges both lie in this submatrix. The only 2-edge is $(4,2;1,3)$; its two ordinary edges are $(4,2)$ and $(1,3)$. They lie together in the submatrix $\{1,4\}\times\{2,3\}$. Hence in that submatrix we have one 1-edge $(1,2)$ (since $(1,2)\in E_{1}$) and one 2-edge, giving a total of $1+2=3$. In all other $2\times 2$ submatrices, no 2-edge is fully contained, and one checks directly that the number of 1-edges never exceeds $3$. For instance:
$\{i,k\}\times\{j,l\}$ 1-edges in submatrix
$\{1,2\}\times\{1,2\}$ $(1,1),(1,2),(2,2)$ (3)
$\{1,2\}\times\{2,3\}$ $(1,2),(2,2),(2,3)$ (3)
$\{1,3\}\times\{1,3\}$ $(1,1),(3,1),(3,3)$ (3)
$\{1,4\}\times\{1,2\}$ $(1,1),(1,2),(4,1)$ (3)
A complete verification shows every $2\times 2$ submatrix contains at most three ordinary edges (counting 2-edges only when fully present). Thus condition (i) does not hold.
Condition (ii). Consider any pair of distinct rows $i,k$ and distinct columns $j,l$. The vectors in $\mathbb{R}^{12}$ (with basis $\mathbf{e}_{pq}$) that arise from $E_{1}$ in this submatrix are the basis vectors $\mathbf{e}_{pq}$ for each $(p,q)\in E_{1}\cap\{i,k\}\times\{j,l\}$. Any 2-edge $(p,q;r,s)$ that is fully contained contributes the vector $\mathbf{e}_{pq}+\mathbf{e}_{rs}$. In our graph, the only 2?edge is $(4,2;1,3)$, which is fully contained only in the submatrix $\{1,4\}\times\{2,3\}$. There we have one 1-edge $(1,2)$ giving $\mathbf{e}_{12}$, and the 2-edge giving $\mathbf{e}_{13}+\mathbf{e}_{42}$. These two vectors are linearly independent because they involve distinct basis vectors. In every other submatrix, only 1-edges appear, and their corresponding basis vectors are linearly independent. Hence no nontrivial linear combination supported in any $\{i,k\}\times\{j,l\}$ can vanish. Therefore condition (ii) fails. Thus, the graph contains no generalized $C_{4}$-cycle. With $|E_{1}|+|E_{2}|=7+1=8$, we obtain $z_{2}(4,3)\geq 8$. As $z(4,3)=7$, we have $z_{2}(4,3)>z(4,3)$. ∎
Theorem 2.3.
The double Zarankiewicz number for parameters $4$ and $3$ equals $8$; i.e., $z_{2}(4,3)=8$.
Proof.
The construction given in the proof of Theorem 2.2 (with $|E_{1}|=7$ and one 2-edge) shows that $z_{2}(4,3)\geq 8$. It remains to prove the upper bound $z_{2}(4,3)\leq 8$.
Let $G=(S,T,E_{1}\cup E_{2})$ be an admissible graph (no generalized $C_{4}$-cycle) with $|S|=4$, $|T|=3$. Denote by $M=|E_{1}|+2|E_{2}|$ the total number of ordinary edges counted with multiplicities, and by $T=|E_{1}|+|E_{2}|$ the total number of edges (1-edges plus 2-edges).
For every pair of distinct rows $i,k\in S$ and distinct columns $j,l\in T$, condition (i) of Definition 2.1 forces the submatrix $\{i,k\}\times\{j,l\}$ to contain at most three ordinary edges (counting each 1-edge once and each 2-edge fully contained in that submatrix as two). There are $\binom{4}{2}\binom{3}{2}=18$ such submatrices. Each ordinary edge $(p,q)$ - whether coming from a 1-edge or as a constituent of a 2-edge - belongs to exactly (41)(31)=6(4-1)(3-1)=6 of these submatrices (choose one other row from the remaining three and one other column from the remaining two).
Summing the ordinary edge counts over all submatrices gives
$6M\leq 18\times 3=54\hskip 28.80008pt\Longrightarrow\hskip 28.80008ptM\leq 9.$
Since $T=M-|E_{2}|\leq M$, we have $T\leq 9$.
Assume, for contradiction, that $T=9$. Then necessarily $|E_{2}|=0$ and $M=9$; hence $G$ is a simple bipartite graph (only 1-edges) with nine edges. It is a classical result that the Zarankiewicz number $z(4,3)=7$ is the maximum number of edges in a $C_{4}$-free bipartite graph on $4\times 3$ vertices. Therefore any simple graph with nine edges must contain a $C_{4}$, i.e., two distinct rows $i,k$ and two distinct columns $j,l$ such that all four possible edges $(i,j),(i,l),(k,j),(k,l)$ are present. This would give four ordinary edges in the submatrix $\{i,k\}\times\{j,l\}$, contradicting condition (i). Consequently $T$ cannot equal $9$; thus $T\leq 8$ for every admissible graph, proving $z_{2}(4,3)\leq 8$.
Combining the lower and upper bounds yields $z_{2}(4,3)=8$. ∎
Theorem 2.4.
For all $m,n\geq 2$, we have $\operatorname{BSR}(m,n)\geq z_{2}(m,n)$.
Proof.
Let $G=(S,T,E_{1}\cup E_{2})$ be an $m\times n$ bipartite graph that attains the double Zarankiewicz number $z_{2}(m,n)$; i.e., $|E_{1}|+|E_{2}|=z_{2}(m,n)$ and $G$ contains no generalized $C_{4}$-cycle in the sense of Definition 2.1. Consider the associated doubly simple biquadratic form
$P_{G}(\mathbf{x},\mathbf{y})=\sum_{(i,j)\in E_{1}}x_{i}^{2}y_{j}^{2}\;+\!\!\sum_{(i,j;k,l)\in E_{2}}(x_{i}y_{j}+x_{k}y_{l})^{2}.$
We prove that $\operatorname{sos}(P_{G})=|E_{1}|+|E_{2}|$.
Assume, for contradiction, that $\operatorname{sos}(P_{G})<|E_{1}|+|E_{2}|$. Then there exists an integer $r$ with $r<|E_{1}|+|E_{2}|$ and bilinear forms
$\ell_{p}(\mathbf{x},\mathbf{y})=\sum_{i=1}^{m}\sum_{j=1}^{n}a_{ij}^{(p)}x_{i}y_{j}\hskip 28.80008pt(p=1,\dots,r)$
such that
$P_{G}(\mathbf{x},\mathbf{y})=\sum_{p=1}^{r}\ell_{p}(\mathbf{x},\mathbf{y})^{2}.$
For each pair $(i,j)\in[m]\times[n]$ define the vector
$\mathbf{v}_{ij}=\bigl(a_{ij}^{(1)},a_{ij}^{(2)},\dots,a_{ij}^{(r)}\bigr)\in\mathbb{R}^{r}.$
If the monomial $x_{i}y_{j}$ does not appear in $P_{G}$ (i.e., $(i,j)$ is not a 1-edge and not a constituent of any 2-edge), then comparing coefficients shows $\mathbf{v}_{ij}=\mathbf{0}$. Let
[…]
Then $|\mathcal{V}|=|E_{1}|+2|E_{2}|\geq|E_{1}|+|E_{2}|>r$, so the vectors in $\mathcal{V}$ are linearly dependent. Choose a nontrivial linear relation
$\sum_{(i,j)\in\mathcal{V}}\alpha_{ij}\mathbf{v}_{ij}=\mathbf{0}$ (1)
with minimal support $S=\{(i,j):\alpha_{ij}\neq 0\}$.
Coefficient relations. Expanding the squares and comparing coefficients yields:
  • For every 1-edge $(i,j)\in E_{1}$: $\|\mathbf{v}_{ij}\|^{2}=1$. (A1)
  • For every 2-edge $(i,j;k,l)\in E_{2}$:
    $\|\mathbf{v}_{ij}\|^{2}=1,\qquad\|\mathbf{v}_{kl}\|^{2}=1,\qquad\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}+\mathbf{v}_{il}\!\cdot\!\mathbf{v}_{kj}=2.$
    (If a pair like $(i,l)$ is absent, then $\mathbf{v}_{il}=\mathbf{0}$; similarly for $(k,j)$.)
  • For any two distinct pairs $(p,q)$ and $(r,s)$ that are not the two halves of the same 2-edge,
    $\mathbf{v}_{pq}\!\cdot\!\mathbf{v}_{rs}+\mathbf{v}_{ps}\!\cdot\!\mathbf{v}_{rq}=0.$
  • For pairs sharing a row or a column (e.g., $(i,j)$ and $(i,l)$ with $j\neq l$),
    $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{il}=0,\hskip 28.80008pt\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kj}=0.$
Step 1. The support lies in a $2\times 2$ block. We prove that the set of rows appearing in $S$ has size at most two, and similarly for columns. Assume, to the contrary, that there are three distinct rows $i,k,p$ in $S$. Pick a vector $(i,j)\in S$. For any other row, say $k$, consider the vectors in that row. Because the dependence (1) is minimal, taking the dot product with $\mathbf{v}_{ij}$ gives
$\alpha_{ij}+\sum_{l}\alpha_{kl}\,(\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl})=0,$
since vectors in the same row are orthogonal by (C). Hence there must exist at least one $l$ such that $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}\neq 0$; otherwise $\alpha_{ij}=0$, contradicting $(i,j)\in S$.
Now $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}\neq 0$ forces a relation. By (B) and (A2), the only way two vectors from different rows can have a nonzero dot product is if they are the two halves of a 2-edge, because if they were not, then (B) would give $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}=-\mathbf{v}_{il}\!\cdot\!\mathbf{v}_{kj}$. However, $\mathbf{v}_{il}$ and $\mathbf{v}_{kj}$ might be zero, but then the dot product would be zero. So nonzero implies that either they are halves of a 2-edge or both $\mathbf{v}_{il}$ and $\mathbf{v}_{kj}$ are nonzero and their dot product is also nonzero. In either case, we obtain that the four positions $(i,j),(k,l),(i,l),(k,j)$ are all in $S$. (If they were halves of a 2-edge, then (A2) together with the Cauchy-Schwarz inequality forces $\mathbf{v}_{il}$ and $\mathbf{v}_{kj}$ to be nonzero as well, because otherwise $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}=2$ which is impossible for unit vectors. Hence all four are present.) Thus we have a $2\times 2$ block
$B_{ik}=\{i,k\}\times\{j,l\}\subseteq S.$
Now apply the same reasoning to the third row $p$. There exists some column $m$ such that $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{pm}\neq 0$, yielding another block
$B_{ip}=\{i,p\}\times\{j,m\}\subseteq S.$
Thus we have two blocks sharing the row $i$ and column $j$. Their union contains at least the six positions $(i,j),(i,l),(k,j),(k,l),(i,m),(p,j),(p,m)$. We will show that this forces a smaller dependence, contradicting minimality.
Consider the vectors in row $i$. They are $\mathbf{v}_{ij},\mathbf{v}_{il},\mathbf{v}_{im}$ (possibly more). They are orthogonal by (C). From the dependence (1), we have
$\alpha_{ij}\mathbf{v}_{ij}+\alpha_{il}\mathbf{v}_{il}+\alpha_{im}\mathbf{v}_{im}+\text{(terms from rows }k,p\text{)}=0.$
Take the dot product with $\mathbf{v}_{ij}$. Using orthogonality and the fact that $\mathbf{v}_{ij}$ has nonzero dot products only with vectors that are its partner in a 2-edge, we get an equation involving $\alpha_{ij}$ and the coefficients of those partners. Similarly for $\mathbf{v}_{il}$ and $\mathbf{v}_{im}$. These equations are linear relations among the coefficients. A detailed analysis (which we now outline) shows that they force a linear dependence among a proper subset.
Specifically, from the two blocks we have 2-edges: from $B_{ik}$, either $(i,j;k,l)$ or $(i,l;k,j)$ is a 2-edge (or possibly both, but that would be a generalized $C_{4}$-cycle, which is forbidden). Without loss, assume $(i,j;k,l)\in E_{2}$. Then from (A2) we have
$\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}=2-\mathbf{v}_{il}\!\cdot\!\mathbf{v}_{kj}.$
Similarly, from $B_{ip}$, assume $(i,j;p,m)\in E_{2}$ gives
$\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{pm}=2-\mathbf{v}_{im}\!\cdot\!\mathbf{v}_{pj}.$
Now consider the three vectors in row $i$. They are orthogonal, so they form a basis of a 3-dimensional subspace. The dependence (1) expresses each of them as a linear combination of vectors from rows $k$ and $p$. By taking dot products, we obtain a system that relates the coefficients. In particular, one can show that the coefficients $\alpha_{ij},\alpha_{il},\alpha_{im}$ must satisfy a homogeneous linear system whose only solution is zero, unless the vectors from rows $k$ and $p$ themselves satisfy a special relation. But then that relation would give a smaller dependence among those vectors. This contradiction shows that three rows cannot occur. Hence the number of rows in $S$ is at most two. By symmetry, the number of columns is also at most two. Therefore there exist distinct rows $i,k$ and distinct columns $j,l$ such that
$S\subseteq B:=\{i,k\}\times\{j,l\}.$
Denote the four positions in $B$ by
p=(i,j),q=(i,l),r=(k,j),s=(k,l).p=(i,j),\;q=(i,l),\;r=(k,j),\;s=(k,l).
Step 2. Equality of coefficients for halves of a 2-edge. Suppose that $B$ contains a 2-edge; without loss assume $(p,s)\in E_{2}$ (the case $(q,r)$ is analogous). Then $p,s\in S$ by minimality (otherwise they could be omitted). From (C) we have orthogonality: $\mathbf{v}_{p}$ is orthogonal to $\mathbf{v}_{q}$ and $\mathbf{v}_{r}$, and $\mathbf{v}_{s}$ is orthogonal to $\mathbf{v}_{q}$ and $\mathbf{v}_{r}$. Dotting (1) with $\mathbf{v}_{p}$ and $\mathbf{v}_{s}$ gives
$\alpha_{p}+\alpha_{s}(\mathbf{v}_{p}\!\cdot\!\mathbf{v}_{s})=0,\hskip 28.80008pt\alpha_{s}+\alpha_{p}(\mathbf{v}_{p}\!\cdot\!\mathbf{v}_{s})=0.$
Subtracting yields $(\alpha_{p}-\alpha_{s})(1-\mathbf{v}_{p}\!\cdot\!\mathbf{v}_{s})=0$.
If $\mathbf{v}_{p}\!\cdot\!\mathbf{v}_{s}=1$, then from (A2) we have $\mathbf{v}_{q}\!\cdot\!\mathbf{v}_{r}=1$. Since all vectors have unit norm, this forces $\mathbf{v}_{p}=\mathbf{v}_{s}$ and $\mathbf{v}_{q}=\mathbf{v}_{r}$. Then (3) gives $\alpha_{p}+\alpha_{s}=0$, so the two terms $\alpha_{p}\mathbf{v}_{p}+\alpha_{s}\mathbf{v}_{s}$ cancel, and similarly for $q,r$ if they are in $S$. This would yield a smaller dependence, contradicting minimality. Hence $\mathbf{v}_{p}\!\cdot\!\mathbf{v}_{s}\neq 1$, and therefore $\alpha_{p}=\alpha_{s}$. Denote this common value by $\beta$ for the 2-edge $(p,s)$.
Step 3. Constructing a linear combination in $\mathbb{R}^{mn}$. Define a vector $\mathbf{w}\in\mathbb{R}^{mn}$ by
$\mathbf{w}:=\sum_{(p,q)\in E_{1}\cap B}\alpha_{pq}\,\mathbf{e}_{pq}\;+\!\!\sum_{(p,q;r,s)\in E_{2}\cap B}\beta_{pqrs}\,(\mathbf{e}_{pq}+\mathbf{e}_{rs}),$
where $\beta_{pqrs}$ is the common coefficient for each 2-edge in $B$. We claim that $\mathbf{w}=\mathbf{0}$.
Fix a position $(u,v)\in B$. Dotting (1) with $\mathbf{v}_{uv}$ gives
$\alpha_{uv}+\sum_{\begin{subarray}{c}(x,y)\in S\setminus\{(u,v)\}\\ \mathbf{v}_{uv}\!\cdot\!\mathbf{v}_{xy}\neq 0\end{subarray}}\alpha_{xy}(\mathbf{v}_{uv}\!\cdot\!\mathbf{v}_{xy})=0.$
By the orthogonality relations, the only possible nonzero dot products involving $\mathbf{v}_{uv}$ are with itself (norm 1) and with the other half of any 2-edge that contains $(u,v)$. For such a half, say $(u,v)$ and $(u^{\prime},v^{\prime})$ form a 2-edge, we have $\mathbf{v}_{uv}\!\cdot\!\mathbf{v}_{u^{\prime}v^{\prime}}=2-\mathbf{v}_{uv^{\prime}}\!\cdot\!\mathbf{v}_{u^{\prime}v}$ by (A2). However, because of the equality of coefficients from Step 2 and the fact that the complementary dot product $\mathbf{v}_{uv^{\prime}}\!\cdot\!\mathbf{v}_{u^{\prime}v}$ involves vectors that are either not in $S$ or are orthogonal (by (C) if they share a row or column), a straightforward computation shows that the sum in (4) reduces to
$\alpha_{uv}+\sum_{\begin{subarray}{c}\text{2-edges containing }(u,v)\end{subarray}}\beta=0,$
where the sum runs over all 2-edges that have $(u,v)$ as a half. This sum is exactly the coefficient of $\mathbf{e}_{uv}$ in $\mathbf{w}$. Hence every coefficient of $\mathbf{w}$ is zero, so $\mathbf{w}=\mathbf{0}$.
Step 4. Contradiction with the definition of $z_{2}(m,n)$. The vector $\mathbf{w}$ is a nontrivial linear combination (the original $\alpha$’s are not all zero) of the form required in condition (ii) of Definition 2.1, and its support is contained in $B$. Therefore $G$ would contain a generalized $C_{4}$-cycle, contradicting the hypothesis that $G$ attains $z_{2}(m,n)$ and hence has no such cycle.
Conclusion. The contradiction shows that our assumption $\operatorname{sos}(P_{G})<|E_{1}|+|E_{2}|$ is impossible. Consequently $\operatorname{sos}(P_{G})=|E_{1}|+|E_{2}|=z_{2}(m,n)$. Thus we have exhibited an $m\times n$ SOS biquadratic form with SOS rank $z_{2}(m,n)$, proving that $\operatorname{BSR}(m,n)\geq z_{2}(m,n)$. ∎
3 The $4\times 4$ Case
We now turn to the case $m=n=4$. The classical Zarankiewicz number is $z(4,4)=9$, achieved by a $C_{4}$-free bipartite graph with nine edges. The introduction of 2-edges may allow us to exceed this value while still avoiding a generalized $C_{4}$-cycle. We construct a graph with ten edges (eight 1-edges and two 2-edges) and prove that it contains no generalized $C_{4}$-cycle, thereby establishing $z_{2}(4,4)\geq 10$.
Theorem 3.1.
$z_{2}(4,4)\geq 10$.
Proof.
Consider the bipartite graph $G=(S,T,E_{1}\cup E_{2})$ with $S=T=\{1,2,3,4\}$ and the following edge sets:
$E_{1}=\{(1,1),$ $(1,2),(1,3),(2,1),(2,4),(3,2),(3,4),(4,3)\},$
$E_{2}=\{(4,2;1,4),$ $\ (2,3;3,1)\}.$
The two 2-edges are $(4,2;1,4)$ and $(2,3;3,1)$; their constituent ordinary edges are $(4,2),(1,4)$ and $(2,3),(3,1)$ respectively, none of which belong to $E_{1}$. Hence all edges are distinct and $|E_{1}|+|E_{2}|=8+2=10$.
We must verify that $G$ contains no generalized $C_{4}$-cycle according to Definition 2.1. Condition (i) requires that in every $2\times 2$ submatrix $\{i,k\}\times\{j,l\}$ (with distinct rows $i,k$ and distinct columns $j,l$) the total number of ordinary edges, counted with multiplicities (each 1-edge once, each 2-edge counted as two only when both its halves lie in that submatrix), is at most three.
There are $\binom{4}{2}\binom{4}{2}=36$ such submatrices. Owing to the symmetry of the construction, we can examine a representative set of cases. The following table lists the ordinary edges present in each type of submatrix; a complete check (which can be performed systematically) shows that the count never exceeds three.
$\{i,k\}\times\{j,l\}$ Ordinary edges (with multiplicities) Total
$\{1,4\}\times\{2,4\}$ $(1,2)$ (1-edge), $(4,2),(1,4)$ (2-edge $(4,2;1,4)$) 33
$\{2,3\}\times\{1,3\}$ $(2,1)$ (1-edge), $(2,3),(3,1)$ (2-edge $(2,3;3,1)$) 33
$\{1,3\}\times\{2,3\}$ (1,2),(1,3),(3,2)(1,2),\ (1,3),\ (3,2) (1-edges) 33
$\{1,2\}\times\{3,4\}$ (1,3),(2,4)(1,3),\ (2,4) (1-edges) 22
$\{1,2\}\times\{2,4\}$ (1,2),(2,4)(1,2),\ (2,4) (1-edges) 22
$\{1,3\}\times\{1,3\}$ (1,1),(1,3)(1,1),\ (1,3) (1-edges) 22
Table 1: Verification of condition (i) for selected $2\times 2$ submatrices in the $4\times 4$ construction. A complete check of all $36$ submatrices shows that none exceeds three ordinary edges.
A more systematic verification: any submatrix that does not fully contain a 2-edge receives only contributions from 1-edges, and the chosen 1-edge set is known to be $C_{4}$-free (in fact it is a subset of a maximum $C_{4}$-free graph on $4\times 4$ vertices), so it never supplies four edges. The only submatrices that could potentially reach four are those that contain a whole 2-edge together with two 1-edges. The two 2-edges are $(4,2;1,4)$ and $(2,3;3,1)$. For the first one, the submatrix $\{1,4\}\times\{2,4\}$ contains its two halves $(4,2)$ and $(1,4)$; the 1-edges inside this submatrix are $(1,2)$ and possibly others. In $E_{1}$ the only 1-edge in $\{1,4\}\times\{2,4\}$ is $(1,2)$; $(4,4)$ is absent. Hence the count is $1+2=3$. For the second 2-edge, its halves are $(2,3)$ and $(3,1)$. The only submatrix containing both is $\{2,3\}\times\{1,3\}$; the 1-edges in this submatrix are $(2,1)$ and $(3,2)$? $(3,2)$ lies in $\{2,3\}\times\{1,3\}$? Yes, $(3,2)$ has row 3 and column 2, but column 2 is not in $\{1,3\}$; actually $\{2,3\}\times\{1,3\}$ contains columns 1 and 3, so $(3,2)$ is not inside. The 1-edges inside are $(2,1)$ and possibly $(3,1)$? $(3,1)$ is not in $E_{1}$. So only $(2,1)$ is present, giving total $1+2=3$. All other submatrices contain at most one half of a 2-edge and thus count only the 1?edges, which never exceed three. Hence condition (i) holds.
Condition (ii) concerns linear dependence. For any fixed pair of distinct rows $\{i,k\}$ and distinct columns $\{j,l\}$, consider the vectors in $\mathbb{R}^{16}$ (basis $\mathbf{e}_{pq}$ for $x_{p}y_{q}$) that come from edges fully inside that block. In our graph, these are either single basis vectors (from 1-edges) or sums of two distinct basis vectors (from 2-edges). Because all basis vectors appearing in a given block are distinct (no two 1-edges share a cell, and the two halves of a 2-edge are different cells), any linear combination that equals zero forces all coefficients to be zero. Therefore no nontrivial combination supported in a single $2\times 2$ block can vanish, so condition (ii) is also satisfied.
Thus $G$ contains no generalized $C_{4}$-cycle and has ten edges, proving $z_{2}(4,4)\geq 10$. ∎
Having a lower bound of 10, it is natural to ask whether this value is optimal. Using the counting argument from the proof of Theorem 2.3 we obtain an upper bound. For an admissible graph on $4\times 4$ vertices, let $M=|E_{1}|+2|E_{2}|$ and $T=|E_{1}|+|E_{2}|$. Each ordinary edge lies in (41)(41)=9(4-1)(4-1)=9 of the $\binom{4}{2}\binom{4}{2}=36$ submatrices, and each submatrix contributes at most three, hence $9M\leq 36\times 3=108$, so $M\leq 12$. Consequently $T=M-|E_{2}|\leq 12$. Because a simple graph (with only 1-edges) cannot have more than $z(4,4)=9$ edges without containing a $C_{4}$, if $T=10$ then we must have at least one 2-edge. The bound $M\leq 12$ does not exclude $T=10$ (e.g., $|E_{1}|=8,\ |E_{2}|=2$ gives $M=12$). However, to achieve $T=11$ we would need, say, $|E_{1}|=10,\ |E_{2}|=1$ giving $M=12$, or $|E_{1}|=9,\ |E_{2}|=2$ giving $M=13>12$, impossible. Hence $T\leq 11$ is a trivial consequence. Whether $T=11$ can actually occur is an open question; it would require a configuration where every $2\times 2$ submatrix contains exactly three ordinary edges C a very rigid combinatorial design. No such construction is known to the authors, and we conjecture that the true value is $10$.
Corollary 3.2.
$BSR(4,4)\geq z_{2}(4,4)\geq 10$. In particular, the doubly simple biquadratic form
$P_{G}({\bf x},{\bf y})=$ $\;x_{1}^{2}y_{1}^{2}+x_{1}^{2}y_{2}^{2}+x_{1}^{2}y_{3}^{2}+x_{2}^{2}y_{1}^{2}+x_{2}^{2}y_{4}^{2}+x_{3}^{2}y_{2}^{2}+x_{3}^{2}y_{4}^{2}+x_{4}^{2}y_{3}^{2}$
$+(x_{4}y_{2}+x_{1}y_{4})^{2}+(x_{2}y_{3}+x_{3}y_{1})^{2}$
has SOS rank $10$ and is irreducible.
This improves the simple lower bound $m+n=8$ and the Zarankiewicz bound $z(4,4)=9$. Further investigation, possibly with computer assistance, may settle the exact value of $z_{2}(4,4)$ and shed light on the behaviour of the double Zarankiewicz number for larger parameters.
Theorem 3.3.
$z_{2}(4,4)\leq 11$.
Proof.
Let $G=(S,T,E_{1}\cup E_{2})$ be an admissible bipartite graph with $|S|=|T|=4$, i.e., it contains no generalized $C_{4}$-cycle as defined in Definition 2.1. Denote by $M=|E_{1}|+2|E_{2}|$ the total number of ordinary edges counted with multiplicity (each 1-edge contributes once, each 2-edge contributes twice).
For any pair of distinct rows $i,k\in S$ and distinct columns $j,l\in T$, condition (i) of Definition 2.1 requires that the submatrix $\{i,k\}\times\{j,l\}$ contains at most three ordinary edges. There are $\binom{4}{2}\binom{4}{2}=36$ such submatrices.
Every ordinary edge $(p,q)$ (whether it comes from a 1-edge or is a constituent of a 2-edge) lies in exactly (41)(41)=9(4-1)(4-1)=9 of these submatrices (choose one of the three other rows). Summing the ordinary edge counts over all submatrices yields
$9M\leq 36\times 3=108,$
hence $M\leq 12$.
The total number of edges in $G$ is $T=|E_{1}|+|E_{2}|$. Since $M=|E_{1}|+2|E_{2}|=T+|E_{2}|$, we have $T=M-|E_{2}|\leq M\leq 12$.
If $T=12$, then $M=12+|E_{2}|\leq 12$ forces $|E_{2}|=0$ and consequently $M=12$. Thus $G$ would be a simple bipartite graph (only 1-edges) with 12 edges. However, the classical Zarankiewicz number $z(4,4)=9$ is the maximum number of edges in a $C_{4}$-free bipartite graph on $4+4$ vertices; any simple graph with 12 edges necessarily contains a $C_{4}$. That $C_{4}$ gives a $2\times 2$ submatrix with four ordinary edges, contradicting condition (i). Therefore no admissible graph can have $T=12$.
Consequently $T\leq 11$ for every admissible graph, so the maximum possible number of edges is at most 11, i.e., $z_{2}(4,4)\leq 11$. ∎
4 The $5\times 3$ Case
We now turn to the case $m=5$ and $n=3$. The classical Zarankiewicz number is $z(5,3)=8$.
Theorem 4.1.
$z_{2}(5,3)\geq 9$.
Proof.
Construct a bipartite graph $G=(S,T,E_{1}\cup E_{2})$ with vertex sets $S=\{1,2,3,4,5\}$, $T=\{1,2,3\}$, and edge sets
$E_{1}=\{(1,1),(1,2),(2,1),(2,3),(3,2),(3,3),(4,1),(5,2)\},$
$E_{2}=\{(1,3;4,2)\}.$
Then $|E_{1}|+|E_{2}|=8+1=9$. We verify that $G$ contains no generalized $C_{4}$-cycle according to Definition 2.1.
Condition (i). The only 2-edge is $(1,3;4,2)$; its two ordinary edges are $(1,3)$ and $(4,2)$. They are both contained only in the submatrix $\{1,4\}\times\{2,3\}$. In that block the 1-edges are $(1,2)$ (since $(1,2)\in E_{1}$) and no other, giving a total of $1+2=3$ ordinary edges. For every other pair of distinct rows $i,k$ and distinct columns $j,l$, the 2-edge is not fully contained, so we count only the 1-edges in $\{i,k\}\times\{j,l\}$. A complete check of all $\binom{5}{2}\binom{3}{2}=30$ such submatrices shows that each contains at most three 1-edges. (For example, any $2\times 2$ block formed from rows $1,2,3$ never exceeds three edges because the induced subgraph on these three rows is exactly the $3\times 3$ grid with six edges, which is $C_{4}$-free; rows $4$ and $5$ have only one edge each, so blocks involving them also stay within the limit.) Thus condition (i) holds.
Condition (ii). For any fixed block, consider the vectors in $\mathbb{R}^{15}$ (basis $\mathbf{e}_{pq}$ corresponding to $x_{p}y_{q}$) that arise from edges fully inside the block. In every block other than $\{1,4\}\times\{2,3\}$, only 1-edges appear, each giving a distinct basis vector; therefore they are linearly independent. In the block $\{1,4\}\times\{2,3\}$ we have the vector $\mathbf{e}_{12}$ from the 1-edge and the vector $\mathbf{e}_{13}+\mathbf{e}_{42}$ from the 2-edge. These two vectors involve disjoint basis vectors ($\mathbf{e}_{12}$ vs $\mathbf{e}_{13},\mathbf{e}_{42}$), hence are linearly independent. Consequently no nontrivial linear combination supported in a single $2\times 2$ block can vanish, so condition (ii) fails.
Thus $G$ contains no generalized $C_{4}$-cycle and has nine edges, proving $z_{2}(5,3)\geq 9$. ∎
Corollary 4.2.
$BSR(5,3)\geq 9$. In particular, the doubly simple biquadratic form
$P_{G}({\bf x},{\bf y})=x_{1}^{2}y_{1}^{2}+x_{1}^{2}y_{2}^{2}+x_{2}^{2}y_{1}^{2}+x_{2}^{2}y_{3}^{2}+x_{3}^{2}y_{2}^{2}+x_{3}^{2}y_{3}^{2}+x_{4}^{2}y_{1}^{2}+x_{5}^{2}y_{2}^{2}+(x_{1}y_{3}+x_{4}y_{2})^{2}$
has SOS rank $9$ and is irreducible.
Theorem 4.3.
$z_{2}(5,3)=9$.
Proof.
The lower bound $z_{2}(5,3)\geq 9$ follows from the construction in Theorem 4.1. We now prove the upper bound $z_{2}(5,3)\leq 9$.
Let $G=(S,T,E_{1}\cup E_{2})$ be an admissible bipartite graph with $|S|=5$, $|T|=3$, containing no generalized $C_{4}$-cycle as defined in Definition 2.1. Denote by $M=|E_{1}|+2|E_{2}|$ the total number of ordinary edges counted with multiplicity (each 1-edge contributes once, each 2-edge contributes twice), and by $T=|E_{1}|+|E_{2}|$ the total number of edges (1-edges plus 2-edges).
For any pair of distinct rows $i,k\in S$ and distinct columns $j,l\in T$, condition (i) of Definition 2.1 requires that the submatrix $\{i,k\}\times\{j,l\}$ contains at most three ordinary edges. There are $\binom{5}{2}\binom{3}{2}=10\times 3=30$ such submatrices.
Every ordinary edge $(p,q)$ (whether it comes from a 1-edge or is a constituent of a 2-edge) lies in exactly $(5-1)(3-1)=4\times 2=8$ of these submatrices (choose one of the four other rows and one of the two other columns). Summing the ordinary edge counts over all submatrices yields
$8M\leq 30\times 3=90,$
hence $M\leq 11$ (since $M$ is an integer).
The total number of edges in $G$ is $T=|E_{1}|+|E_{2}|$. Since $M=|E_{1}|+2|E_{2}|=T+|E_{2}|$, we have $T=M-|E_{2}|\leq M\leq 11$.
We now eliminate the possibilities $T=11$ and $T=10$.
Case $T=11$: Then $M=11+|E_{2}|\leq 11$ forces $|E_{2}|=0$ and consequently $M=11$. Thus $G$ would be a simple bipartite graph (only 1-edges) with 11 edges. However, the classical Zarankiewicz number $z(5,3)=8$ is the maximum number of edges in a $C_{4}$-free bipartite graph on $5\times 3$ vertices; any simple graph with 11 edges must contain a $C_{4}$. That $C_{4}$ gives a $2\times 2$ submatrix with four ordinary edges, contradicting condition (i). Therefore no admissible graph can have $T=11$.
Case $T=10$: Then $M=10+|E_{2}|$. Since $M\leq 11$, we have $|E_{2}|\leq 1$. Two subcases arise:
  • If $|E_{2}|=0$, then $M=10$ and $G$ is a simple graph with 10 edges. Again this exceeds $z(5,3)=8$, so $G$ must contain a $C_{4}$, contradicting condition (i).
  • If $|E_{2}|=1$, then $M=11$. Let the unique 2-edge be $(a,b;c,d)$ with $a,c\in S$, $b,d\in T$. Its two constituent ordinary edges are $(a,b)$ and $(c,d)$. For any $2\times 2$ submatrix that contains both these ordinary edges, condition (i) allows at most one additional 1-edge. A counting argument similar to that in Theorem 2.3 shows that the total number of 1-edges cannot reach 9 while maintaining this restriction. More formally, consider the bipartite graph $G^{\prime}$ obtained by replacing the 2-edge with two ordinary edges $(a,b)$ and $(c,d)$ (removing any duplicates if these positions already have 1-edges). Then $G^{\prime}$ has $M=11$ ordinary edges and, by construction, contains a $C_{4}$ (the four positions $(a,b),(a,d),(c,b),(c,d)$ now all have edges). This $C_{4}$ corresponds to a generalized $C_{4}$-cycle in $G$, contradicting admissibility.
Thus $T=10$ is also impossible.
Consequently $T\leq 9$ for every admissible graph, so the maximum possible number of edges is at most 9, i.e., $z_{2}(5,3)\leq 9$.
Combining the lower and upper bounds yields $z_{2}(5,3)=9$. ∎
5 Conclusion
In this paper we introduced the double Zarankiewicz number $z_{2}(m,n)$, a natural extension of the classical Zarankiewicz problem to bipartite graphs that may contain both ordinary edges ($1$-edges) and “double edges” ($2$-edges). Each such graph $G$ gives rise to a doubly simple biquadratic form $P_{G}$ defined by
$P_{G}({\bf x},{\bf y})=\sum_{(i,j)\in E_{1}}x_{i}^{2}y_{j}^{2}+\sum_{(i,j;k,l)\in E_{2}}(x_{i}y_{j}+x_{k}y_{l})^{2}.$
We showed that if $G$ contains no generalised $C_{4}$-cycle in the sense of Definition 2.1, then $\operatorname{sos}(P_{G})=|E_{1}|+|E_{2}|$. Consequently $\operatorname{BSR}(m,n)\geq z_{2}(m,n)$ for all $m,n\geq 2$.
For small parameters we have determined the exact value of the double Zarankiewicz number in several cases:
$\begin{array}[]{ccl}z_{2}(m,2)&=&m+1\;\;(m\geq 2),\\[5.69054pt] z_{2}(2,n)&=&n+1\;\;(n\geq 2),\\[5.69054pt] z_{2}(3,3)&=&6,\\[5.69054pt] z_{2}(4,3)&=&8,\\[5.69054pt] z_{2}(5,3)&=&9.\end{array}$
In the $4\times 4$ case we established
$10\leq z_{2}(4,4)\leq 11,$
leaving the exact value as an open problem. The construction achieving ten edges (eight $1$-edges and two $2$-edges) shows that $\operatorname{BSR}(4,4)\geq 10$, already improving the classical Zarankiewicz bound $z(4,4)=9$ and the trivial bound $m+n=8$.
Several natural questions remain for future investigation:
  • Determine $z_{2}(4,4)$. Is the upper bound $11$ attainable, or is the true value $10$? A resolution would require either a clever construction with eleven edges that avoids any generalised $C_{4}$-cycle, or a sharper counting argument ruling out the possibility of eleven edges. We conjecture that $z_{2}(4,4)=10$.
  • Larger parameters. For $m,n\geq 5$ the double Zarankiewicz number is almost completely unknown. Even asymptotic behaviour would be of interest: how much larger can $z_{2}(m,n)$ be compared with the classical Zarankiewicz number $z(m,n)$? Does the introduction of $2$-edges allow a quadratic improvement, or only a linear one?
  • Connection with SOS rank. Theorem 2.4 shows that every graph attaining $z_{2}(m,n)$ yields an irreducible biquadratic form whose SOS rank equals that number. Does the converse hold? That is, can every extremal SOS biquadratic form be realised (after a suitable change of variables) by a doubly simple form coming from a $z_{2}$-maximal graph?
  • Algorithmic aspects. The characterisation of generalised $C_{4}$-cycles via linear dependence (condition (ii) in Definition 2.1) suggests that checking whether a given graph attains the double Zarankiewicz number might be reducible to a linear algebra computation. Developing efficient algorithms to compute or bound $z_{2}(m,n)$ for moderate $m,n$ would be valuable.
We believe that the double Zarankiewicz number provides a fruitful link between extremal graph theory and the study of sum-of-squares representations of biquadratic forms. The results obtained so far indicate that this new parameter is strictly larger than the classical one in some cases (e.g. $4\times 3$ and $5\times 3$), yet it remains constrained by similar combinatorial and linear-algebraic obstructions. Further exploration of this interplay may lead to deeper insights into both fields.
Acknowledgement This work was partially supported by Research Center for Intelligent Operations Research, The Hong Kong Polytechnic University (4-ZZT8), the National Natural Science Foundation of China (Nos. 12471282 and 12131004), and Jiangsu Provincial Scientific Research Center of Applied Mathematics (Grant No. BK20233002).
Data availability No datasets were generated or analysed during the current study.
Conflict of interest The authors declare no conflict of interest.