### 16 Mar 2026 v2

# Biquadratic SOS Rank and Double Zarankiewicz Number

Abstract
The maximum sum-of-squares (SOS) rank of an $m\times n$ biquadratic form, denoted $\operatorname{BSR}(m,n)$, has recently been shown to be bounded below by the classical Zarankiewicz number $z(m,n)$, the maximum number of edges in a $C_{4}$-free bipartite graph. However, a gap was discovered for the $4\times 3$ case: $\operatorname{BSR}(4,3)\geq 8>z(4,3)=7$, arising from a construction that adds a square of a bilinear form $(x_{4}y_{2}+x_{1}y_{3})^{2}$ to a $C_{4}$-free simple biquadratic form. This paper introduces the double Zarankiewicz number $z_{2}(m,n)$ to capture such phenomena. We define a new class of bipartite graphs that may contain both ordinary edges (1-edges) and “double edges” (2-edges), the latter corresponding to squares of simple two-term bilinear forms. A graph $G$ of this type yields a doubly simple biquadratic form $P_{G}$, and we prove that if $G$ contains no generalised $C_{4}$-cycle then $\operatorname{sos}(P_{G})=|E_{1}|+|E_{2}|$; consequently $\operatorname{BSR}(m,n)\geq z_{2}(m,n)$. We determine the exact value of $z_{2}(m,n)$ for several small parameters: $z_{2}(m,2)=m+1$, $z_{2}(2,n)=n+1$, $z_{2}(3,3)=6$, $z_{2}(4,3)=8$, $z_{2}(4,4)=10$, and $z_{2}(5,3)=9$. Furthermore, we establish the lower bound $z_{2}(8,7)\geq 39$, which implies $\operatorname{BSR}(8,7)\geq 39$, improving the previous bound $\operatorname{BSR}(8,7)\geq z(8,7)=37$. These results demonstrate that the double Zarankiewicz number can exceed the classical one by at least two, revealing new algebraic-combinatorial phenomena. The paper concludes with a discussion of open questions and connections between extremal graph theory and the study of SOS representations.
Keywords. Biquadratic form, sum of squares, SOS rank, Zarankiewicz number, double Zarankiewicz number, bipartite graph, $C_{4}$-cycle, extremal graph theory.
AMS subject classifications. 14P10, 05C35, 11E25, 15A69, 90C22.
1 Introduction
Denote $\{1,\ldots,m\}$ as $[m]$. Assume that $m,n\geq 2$. Let
$P(\mathbf{x},\mathbf{y})=\sum_{i,k=1}^{m}\sum_{j,l=1}^{n}a_{ijkl}x_{i}x_{k}y_{j}y_{l}.$ (1)
We call $P$ an $m\times n$ biquadratic form. Here $a_{ijkl}$ are real numbers. We assume that
$a_{ijkl}=a_{kjil}=a_{klij}$ (2)
for $i,k\in[m],j,l\in[n]$. A PSD (positive semi-definite) biquadratic form is one for which $P(\mathbf{x},\mathbf{y})\geq 0$ for all $\mathbf{x},\mathbf{y}$. It is SOS (sum of squares) if it can be written as a finite sum of squares of bilinear forms, i.e.,
$P({\bf x},{\bf y})=\sum_{p=1}^{r}f_{p}({\bf x},{\bf y})^{2},$ (3)
where $f_{p}$ are $m\times n$ bilinear forms for $p\in[r]$. The smallest $r$ such that (3) holds is called the SOS rank of $P$. In this case, we say that $\sum_{p=1}^{r}f_{p}({\bf x},{\bf y})^{2}$ is irreducible. The maximum SOS rank of $m\times n$ SOS biquadratic forms is denoted as $BSR(m,n)$. It is known that $BSR(m,2)=m+1$ and $BSR(2,n)=n+1$ [1], $BSR(3,3)=6$ [2] and $BSR(m,n)\leq mn-1$ [10].
Recently, it was proved [5] that
$BSR(m,n)\geq z(m,n),$ (4)
where $z(m,n)$ is the Zarankiewicz number [13]. The Zarankiewicz problem, originating from [13], asks for the maximum number of edges in an $m\times n$ bipartite graph that contains no complete bipartite subgraph $K_{2,2}$ (i.e., no $C_{4}$). This classical extremal problem has been extensively studied; see [6, 3, 11] for surveys and fundamental bounds.
The inequality (4) is based upon the simple biquadratic forms introduced in [10]. A biquadratic form is a simple biquadratic form if it contains only distinct terms of the type $x_{i}^{2}y_{j}^{2}$. Then we may establish a one-to-one relation. Let $G=(S,T,E)$ be a bipartite graph, where $S=\{1,\cdots,m\}$ and $T=\{1,\cdots,n\}$ are the vertex sets and $E$ is the edge set. Then uniquely we have a simple quadratic form
$P_{G}({\bf x},{\bf y})=\sum_{(i,j)\in E}x_{i}^{2}y_{j}^{2}.$ (5)
If $G$ has no $C_{4}$-cycles, then the simple biquadratic form $P_{G}$ is irreducible, i.e., $\mathrm{sos}(P_{G})=|E|$. On the other hand, the Zarankiewicz number $z(m,n)$ is defined as the largest number of edges such that the $m\times n$ bipartite graph $G$ is $C_{4}$-cycle free [13, 11]. Thus, we have (4).
It is known that $BSR(m,2)=z(m,2)=m+1$, $BSR(2,n)=z(2,n)=n+1$ and $BSR(3,3)=z(3,3)=6$. However, it was found in [12] that $BSR(4,3)\geq 8>z(4,3)=7$. The value $z(4,3)=7$ is a classical result in Zarankiewicz theory [11, 6]. This gap was discovered in the following way. First, we have
$P_{4,3,7}({\bf x},{\bf y})=x_{1}^{2}y_{1}^{2}+x_{2}^{2}y_{2}^{2}+x_{3}^{2}y_{3}^{2}+x_{1}^{2}y_{2}^{2}+x_{2}^{2}y_{3}^{2}+x_{3}^{2}y_{1}^{2}+x_{4}^{2}y_{1}^{2}.$
This is a $4\times 3$ simple biquadratic form whose corresponding $4\times 3$ bipartite graph $G$ attains the Zarankiewicz number $z(4,3)=7.$ Then, an $8$-square $4\times 3$ sos biquadratic form $Q$ is formed as
$Q({\bf x},{\bf y})=P_{4,3,7}({\bf x},{\bf y})+(x_{4}y_{2}+x_{1}y_{3})^{2}.$
It was proved that $Q$ is irreducible, i.e., $\mathrm{sos}(Q)=8$. In this way, it was established that
$BSR(4,3)\geq 8>z(4,3)=7.$ (6)
This makes us think further: Why can such a thing happen? What is its meaning in the theory of the Zarankiewicz number? Now, $(x_{4}y_{2}+x_{1}y_{3})^{2}$ is not a single term square. It does not correspond to an edge of the bipartite graph $G$. But can we interpret it as a super edge of $G$? This motivates this paper.
The remainder of this paper is organized as follows. In Section 2, we introduce the concept of a double Zarankiewicz number $z_{2}(m,n)$ and establish its basic properties, including its relation to the classical Zarankiewicz number and exact values for small parameters. We also prove the fundamental inequality $\operatorname{BSR}(m,n)\geq z_{2}(m,n)$ by showing that any graph attaining the double Zarankiewicz number yields an irreducible doubly simple biquadratic form. Section 3 investigates the $4\times 4$ case, where we prove $z_{2}(4,4)=10$. Section 4 treats the $5\times 3$ case, where we prove $z_{2}(5,3)=9$. Section 5 presents a new lower bound for $z_{2}(8,7)$, showing that $z_{2}(8,7)\geq 39=z(8,7)+2$ and consequently $\operatorname{BSR}(8,7)\geq 39$. Finally, Section 6 concludes the paper with a summary of results and a discussion of open problems.
2 Double Zarankiewicz Number
Let $G=(S,T,E)$ be an $m\times n$ bipartite graph, where $S=\{1,\cdots,m\}$ and $T=\{1,\cdots,n\}$ are its vertex sets. We define the edge set $E=E_{1}\cup E_{2}$, where $E_{1}$ is the $1$-edge set, and $E_{2}$ is the $2$-edge set. A $1$-edge $e$ in $E_{1}$ has the form $e=(i,j)$, where $i\in S$ and $j\in T$. On the other hand, a $2$-edge $e$ in $E_{2}$ has the form $(i,j;k,l)$, where $i,k\in S$ and $j,l\in T$. A $2$-edge is required to be nondegenerate, i.e., $i\neq k$ and $j\neq l$. For such a bipartite graph $G$, we associate it with an $m\times n$ sos biquadratic form $P_{G}$, defined as:
$P_{G}({\bf x},{\bf y})=\sum_{(i,j)\in E_{1}}x_{i}^{2}y_{j}^{2}+\sum_{(i,j;k,l)\in E_{2}}(x_{i}y_{j}+x_{k}y_{l})^{2}.$ (7)
We call such a biquadratic form a doubly simple biquadratic form. If $P_{G}$ is irreducible, then the sos rank of $P_{G}$ is $|E|=|E_{1}|+|E_{2}|$.
Let $m=4$ and $n=3$, $E_{1}=\{(1,1),(2,2),(3,3),(1,2),(2,3),(3,1),(4,1)$ and $E_{2}=\{(i,j;k,l)\}$. Then we see if $(i,j;k,l)$ is $(4,2;1,3)$, $P_{G}$ is irreducible. We wish to characterize the combinatorial feature of such $(i,j;k,l)$.
We now refine the notion of a $C_{4}$-cycle to account for the presence of 2-edges.
Definition 2.1 (Double Zarankiewicz Number).
Let $G=(S,T,E_{1}\cup E_{2})$ be an $m\times n$ bipartite graph with vertex sets $S=[m]$, $T=[n]$, where
  • $E_{1}$ is a set of 1-edges of the form $(i,j)$ with $i\in S$, $j\in T$;
  • $E_{2}$ is a set of 2-edges of the form $(i,j;k,l)$ with $i,k\in S$ and $j,l\in T$. A 2-edge is required to be nondegenerate, i.e. $i\neq k$ and $j\neq l$.
We impose the following simplicity condition:
  1. (S)
    No two distinct edges or subedges of $G$ share a cell. We also assume that $|E_{1}|=z(m,n)$.
We say that $G$ contains a generalized $C_{4}$-cycle if either
  1. 1.
    there exists a classical $C_{4}$-cycle, or
  2. 2.
    there exists a 2-edge $(i,j;k,l)\in E_{2}$ such that both opposite cells $(i,l)$ and $(k,j)$ are occupied, or
  3. 3.
    there exist a 2-edge $(i,j;p,q)$ and a distinct cell $(k,l)$ such that the five cells $(k,l),(k,j),(k,q),(i,l),(p,l)$ are occupied.
The double Zarankiewicz number $z_{2}(m,n)$ is the maximum possible total number of edges $|E_{1}|+|E_{2}|$ for which such a generalized $C_{4}$-cycle does not exist.
Remark 2.2.
Like condition 1, conditions 2 and 3 are also natural and combinatorial. Condition 2 ensures that a 2 edge itself does not create a $C_{4}$ cycle. Because $|E_{1}|=z(m,n)$, each half of a 2 edge would form a $C_{4}$ cycle with some distinct occupied cells (otherwise more 1-edges could be added). Condition?3 then guarantees that the two lists of such cells those completing a $C_{4}$ with the first half and those completing a $C_{4}$ with the second half have no intersection. This symmetric condition is exactly what prevents the linear dependence that would otherwise arise.
Theorem 2.3.
For all $m,n\geq 2$, we have $\operatorname{BSR}(m,n)\geq z_{2}(m,n)\geq z(m,n)$.
Proof.
The inequality $z_{2}(m,n)\geq z(m,n)$ is trivial by taking $E_{2}=\emptyset$. Now let $G=(S,T,E_{1}\cup E_{2})$ be an $m\times n$ bipartite graph attaining $z_{2}(m,n)$; thus $|E_{1}|+|E_{2}|=z_{2}(m,n)$ and $G$ contains no generalized $C_{4}$-cycle in the sense of Definition 2.1. Consider the associated doubly simple biquadratic form
$P_{G}(\mathbf{x},\mathbf{y})=\sum_{(i,j)\in E_{1}}x_{i}^{2}y_{j}^{2}+\sum_{(i,j;k,l)\in E_{2}}(x_{i}y_{j}+x_{k}y_{l})^{2}.$
Assume, for contradiction, that $\mathrm{sos}(P_{G})<|E_{1}|+|E_{2}|$. Then there exists an integer $r<|E_{1}|+|E_{2}|$ and bilinear forms
$\ell_{p}(\mathbf{x},\mathbf{y})=\sum_{i=1}^{m}\sum_{j=1}^{n}a_{ij}^{(p)}x_{i}y_{j}\hskip 28.80008pt(p=1,\dots,r)$
such that
$P_{G}(\mathbf{x},\mathbf{y})=\sum_{p=1}^{r}\ell_{p}(\mathbf{x},\mathbf{y})^{2}.$
For each pair $(i,j)$ define the vector $\mathbf{v}_{ij}=(a_{ij}^{(1)},\dots,a_{ij}^{(r)})\in\mathbb{R}^{r}$. If $(i,j)$ is not a 1-edge and not a half of any 2-edge, then comparing coefficients shows $\mathbf{v}_{ij}=\mathbf{0}$. Let
[…]
By the simplicity condition (S), the $2|E_{2}|$ halves are all distinct and none belongs to $E_{1}$; hence
$|\mathcal{V}|=|E_{1}|+2|E_{2}|\geq|E_{1}|+|E_{2}|>r.$
Thus the vectors in $\mathcal{V}$ are linearly dependent. Choose a nontrivial linear relation
$\sum_{(i,j)\in\mathcal{V}}\alpha_{ij}\mathbf{v}_{ij}=\mathbf{0}$
with minimal support $S=\{(i,j):\alpha_{ij}\neq 0\}$. All vectors in this relation are nonzero and, by minimality, they are distinct (if two were equal we could combine them and obtain a smaller support).
Coefficient relations. Expanding the squares and comparing coefficients yields:
  • For every 1-edge $(i,j)\in E_{1}$: $\|\mathbf{v}_{ij}\|^{2}=1$. (A1)
  • For every 2-edge $(i,j;k,l)\in E_{2}$: $\|\mathbf{v}_{ij}\|^{2}=\|\mathbf{v}_{kl}\|^{2}=1$ and $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}+\mathbf{v}_{il}\!\cdot\!\mathbf{v}_{kj}=1$. (A2)
  • For any two distinct pairs $(p,q),(r,s)$ that are not the two halves of the same 2-edge: $\mathbf{v}_{pq}\!\cdot\!\mathbf{v}_{rs}+\mathbf{v}_{ps}\!\cdot\!\mathbf{v}_{rq}=0$. (B)
  • For pairs sharing a row or a column: $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{il}=0$ (same row), $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kj}=0$ (same column). (C)
Step 1. Consequences of the absence of a generalized $C_{4}$-cycle. Take any 2-edge $(i,j;k,l)\in E_{2}$. Because it is nondegenerate, $i\neq k$ and $j\neq l$. Look at the submatrix $\{i,k\}\times\{j,l\}$. Since no generalized $C_{4}$-cycle exists, at most one of the two opposite cells $(i,l)$ and $(k,j)$ can be occupied (otherwise condition 2 of Definition 2.1 would be violated). Hence at least one of $\mathbf{v}_{il},\mathbf{v}_{kj}$ is zero. From (A2) we then obtain $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}=1$. Because $\|\mathbf{v}_{ij}\|=\|\mathbf{v}_{kl}\|=1$, Cauchy-Schwarz forces $\mathbf{v}_{ij}=\mathbf{v}_{kl}$. Thus each 2-edge contributes two equal vectors; denote this common vector by $\mathbf{w}_{ijkl}$.
Step 2. Orthogonality of distinct vectors. Let $(i,j)$ and $(k,l)$ be two distinct positions in the support $\mathcal{S}$. Apply relation (B) to these two positions:
$\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}+\mathbf{v}_{il}\!\cdot\!\mathbf{v}_{kj}=0.$
Suppose, for contradiction, that $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}\neq 0$. Then $\mathbf{v}_{il}\!\cdot\!\mathbf{v}_{kj}\neq 0$, so both $\mathbf{v}_{il}$ and $\mathbf{v}_{kj}$ are nonzero. Hence the cells $(i,l)$ and $(k,j)$ are occupied (either as 1-edges or as halves of 2-edges). Consequently the four cells
(i,j),(k,l),(i,l),(k,j)(i,j),\;(k,l),\;(i,l),\;(k,j)
are all occupied.
If all four were 1-edges, they would form a classical $C_{4}$, violating condition 1 of Definition 2.1. Therefore at least one of them is a half of a 2-edge. Without loss of generality, assume $(i,j)$ is a half of a 2-edge $e=(i,j;p,q)$ with $p\neq i,\;q\neq j$. Then $\mathbf{v}_{ij}=\mathbf{v}_{pq}$ and $(p,q)$ is occupied.
Now consider the pair $(p,q)$ and $(k,l)$. Their dot product is $\mathbf{v}_{pq}\!\cdot\!\mathbf{v}_{kl}=\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}\neq 0$. Apply (B) to $(p,q)$ and $(k,l)$:
$\mathbf{v}_{pq}\!\cdot\!\mathbf{v}_{kl}+\mathbf{v}_{pl}\!\cdot\!\mathbf{v}_{kq}=0,$
so $\mathbf{v}_{pl}\!\cdot\!\mathbf{v}_{kq}\neq 0$; thus $(p,l)$ and $(k,q)$ are occupied.
We examine the possibilities for $p$ and $q$ relative to $k$ and $l$.
  • If $p=k$, then $(p,l)=(k,l)$ and $(p,q)=(k,q)$. But $(k,q)$ and $(k,l)$ share the same row, so by (C) we have $\mathbf{v}_{kq}\!\cdot\!\mathbf{v}_{kl}=0$. Yet $\mathbf{v}_{kq}=\mathbf{v}_{pq}=\mathbf{v}_{ij}$ and $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}\neq 0$, a contradiction.
  • If $q=l$, then $(p,l)$ and $(k,l)$ share the same column, so (C) gives $\mathbf{v}_{pl}\!\cdot\!\mathbf{v}_{kl}=0$ while $\mathbf{v}_{pl}=\mathbf{v}_{ij}$ again contradicts $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}\neq 0$.
Hence we must have $p\neq k$ and $q\neq l$. In this situation the five cells
(k,l),(k,j),(k,q),(i,l),(p,l)(k,l),\;(k,j),\;(k,q),\;(i,l),\;(p,l)
are all distinct (because $i\neq k$, $j\neq l$, $p\neq k$, $q\neq l$). They are occupied: $(k,l)$ and $(i,l)$ from the original four, $(k,j)$ from the original, $(k,q)$ and $(p,l)$ from the new ones. Moreover, $(i,j;p,q)$ is a 2-edge. Thus the configuration satisfies condition 3 of Definition 2.1, which is forbidden. This contradiction shows that our assumption $\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}\neq 0$ is impossible. Therefore
$\mathbf{v}_{ij}\!\cdot\!\mathbf{v}_{kl}=0\hskip 28.80008pt\text{for any two distinct positions }(i,j),(k,l)\in\mathcal{S}.$
Hence all distinct vectors in the support are pairwise orthogonal.
Step 3. Contradiction. The vectors in the support are nonzero (by (A1) and the fact that halves of 2-edges have norm 1) and pairwise orthogonal, so they are linearly independent. But they satisfy the nontrivial linear relation (1), which is impossible. This contradiction shows that our assumption $\mathrm{sos}(P_{G})<|E_{1}|+|E_{2}|$ was false. Hence $\mathrm{sos}(P_{G})=|E_{1}|+|E_{2}|=z_{2}(m,n)$, proving $\mathrm{BSR}(m,n)\geq z_{2}(m,n)$. ∎
Theorem 2.4.
The double Zarankiewicz number satisfies
$z_{2}(m,2)=m+1\quad\text{for all }m\geq 2,$
$z_{2}(2,n)=n+1\quad\text{for all }n\geq 2,$
$z_{2}(3,3)=6,$
$z_{2}(4,3)\geq 8>z(4,3).$
Proof.
First, for $m=2$ or $n=2$, we have the classical Zarankiewicz numbers $z(m,2)=m+1$ and $z(2,n)=n+1$. By Theorem 2.3, $z_{2}(m,n)\geq z(m,n)$. On the other hand, it is known that $\operatorname{BSR}(m,2)=m+1$ and $\operatorname{BSR}(2,n)=n+1$ [1]. Applying Theorem 2.3 again gives $\operatorname{BSR}(m,n)\geq z_{2}(m,n)$; hence $z_{2}(m,2)\leq m+1$ and $z_{2}(2,n)\leq n+1$. Consequently $z_{2}(m,2)=m+1$ for all $m\geq 2$ and $z_{2}(2,n)=n+1$ for all $n\geq 2$.
For $(3,3)$, the Zarankiewicz number is $z(3,3)=6$ and it is known that $\operatorname{BSR}(3,3)=6$ [2]. The same argument yields $z_{2}(3,3)\geq 6$ and $z_{2}(3,3)\leq\operatorname{BSR}(3,3)=6$; therefore $z_{2}(3,3)=6$.
Finally, for $(4,3)$ we exhibit a construction with total edges $8$. Let
$E_{1}=\{(1,1),(2,2),(3,3),(1,2),(2,3),(3,1),(4,1)\},\hskip 28.80008ptE_{2}=\{(4,2;1,3)\}.$
We verify that this graph contains no generalized $C_{4}$-cycle according to Definition 2.1.
Condition 1. The 1-edges alone form a $C_{4}$-free graph because they attain the Zarankiewicz number $z(4,3)=7$; any two rows share at most one column.
Condition 2. For the unique 2-edge $(4,2;1,3)$, its opposite cells are $(4,3)$ and $(1,2)$. Here $(1,2)\in E_{1}$ but $(4,3)$ is unoccupied, so not both opposite cells are occupied.
Condition 3. We must check that there is no cell $(k,l)$ with $k\in\{1,2,3,4\},\;l\in\{1,2,3\}$ such that the five cells
(k,l),(k,2),(k,3),(4,l),(1,l)(k,l),\;(k,2),\;(k,3),\;(4,l),\;(1,l)
are all occupied. Observe that $(k,2)$ and $(k,3)$ are both occupied only for $k=1$ and $k=2$ (since $(3,2)$ and $(4,3)$ are missing). - For $k=1$: $(1,l)$ is occupied for every $l$, but $(4,l)$ is occupied only for $l=1,2$. * $l=1$ gives the cells $(1,1),(1,2),(1,3),(4,1)$ - only four distinct cells. * $l=2$ gives $(1,2),(1,2),(1,3),(4,2)$ - again four distinct cells. - For $k=2$: $(2,l)$ is occupied for $l=2,3$ (since $(2,1)$ is empty). * $l=2$ gives $(2,2),(2,2),(2,3),(4,2),(1,2)$ - four distinct cells. * $l=3$ would require $(4,3)$, which is missing. Thus no choice yields five *distinct* occupied cells. (The definition of a generalized $C_{4}$-cycle implicitly requires the five cells to be distinct, as used in the proof of Theorem 2.3; otherwise the configuration would already be covered by simpler conditions.)
Therefore $G$ contains no generalized $C_{4}$-cycle. With $|E_{1}|+|E_{2}|=7+1=8$, we obtain $z_{2}(4,3)\geq 8$. Since $z(4,3)=7$, we have $z_{2}(4,3)>z(4,3)$.
Theorem 2.5.
The double Zarankiewicz number for parameters $4$ and $3$ equals $8$; i.e., $z_{2}(4,3)=8$.
Proof.
The construction given in the proof of Theorem 2.4 (with $|E_{1}|=7$ and one 2-edge) shows that $z_{2}(4,3)\geq 8$. We now prove the upper bound $z_{2}(4,3)\leq 8$.
Let $G=(S,T,E_{1}\cup E_{2})$ be an admissible graph (no generalized $C_{4}$-cycle) with $|S|=4$, $|T|=3$. Denote $t=|E_{1}|+|E_{2}|$. Suppose, for contradiction, that $t\geq 9$. Let $k=|E_{2}|$. Then $|E_{1}|=t-k$ and the total number of occupied cells is $|E_{1}|+2k=t+k\leq 12$ (since there are only $12$ cells), so $k\leq 12-t$. For $t=9$ we have $k\leq 3$; for $t=10$ we have $k\leq 2$; etc. We treat the possibilities.
Case $k=0$. Then $G$ is a simple bipartite graph with $|E_{1}|=t\geq 9$. The classical Zarankiewicz number $z(4,3)=7$ implies that any simple graph with $9$ edges must contain a $C_{4}$, contradicting condition 1 of Definition 2.1. Hence $k\geq 1$.
Case $k=3$. Then $|E_{1}|=t-3$. For $t=9$ we have $|E_{1}|=6$ and total occupied cells $6+6=12$; all cells are occupied. For any 2-edge, its two opposite cells are then certainly occupied (since every cell is occupied), violating condition 2. Thus $k\neq 3$. (For $t>9$ we would have $k\leq 12-t<3$, so this case cannot occur.)
Case $k=1$. Then $|E_{1}|=t-1$. For $t=9$ we have $|E_{1}|=8$. The subgraph of 1-edges alone already has $8$ edges, which exceeds the Zarankiewicz number $z(4,3)=7$; therefore it must contain a classical $C_{4}$ (four 1-edges forming a $2\times 2$ submatrix). This contradicts admissibility. Hence $k\neq 1$ for $t=9$.
Case $k=2$. Then $|E_{1}|=t-2$. For $t=9$ we have $|E_{1}|=7$. Thus we have a $C_{4}$-free simple graph with exactly $7$ edges. It is well known (and easy to verify) that any such extremal graph on $4\times 3$ is, up to relabeling of rows and columns, uniquely given by
$E_{1}=\{(1,1),(1,2),(2,2),(2,3),(3,1),(3,3),(4,1)\}$
(or an isomorphic copy). (Indeed, no row can have degree $3$ because that would force the other rows to have total degree at most $3$; the only possible degree sequence is $(2,2,2,1)$, and the three rows of degree $2$ must be incident with the three distinct pairs of columns to avoid a $C_{4}$.) Consequently the set of cells not in $E_{1}$ (the complement) is
$C=\{(1,3),\;(2,1),\;(3,2),\;(4,2),\;(4,3)\}.$
The two 2-edges must use four distinct cells from $C$ (by simplicity conditions), leaving one cell of $C$ empty. Denote the empty cell by $E$. For each 2-edge, condition 2 requires that at least one of its two opposite cells equals $E$ (because all other cells are occupied either by 1-edges or by halves of the other 2-edge). Hence $E$ must be an opposite cell of both 2-edges.
We now examine the five possibilities for $E$ and show that none allows a partition of the remaining four cells into two valid 2-edges with this property. Label the cells of $C$ as
$C_{1}=(1,3),\;C_{2}=(2,1),\;C_{3}=(3,2),\;C_{4}=(4,2),\;C_{5}=(4,3).$
A 2-edge is a pair of cells with distinct rows and distinct columns. The valid pairs among these cells are:
$(C_{1},C_{2}),\;(C_{1},C_{3}),\;(C_{1},C_{4}),\;(C_{2},C_{3}),\;(C_{2},C_{4}),\;(C_{2},C_{5}),\;(C_{3},C_{5}).$
(We omit pairs with equal rows or equal columns.)
For each candidate empty cell we list the remaining four cells and try to split them into two disjoint valid pairs such that each pair has the empty cell as one of its opposites. A direct check shows that no such partition exists:
  • $E=C_{1}=(1,3)$: remaining $\{C_{2},C_{3},C_{4},C_{5}\}$. The only disjoint valid pairing is $\{(C_{2},C_{4}),\;(C_{3},C_{5})\}$. Opposites of $(C_{2},C_{4})$ are $(2,2)$ and $(4,1)$; neither is $(1,3)$.
  • $E=C_{2}=(2,1)$: remaining $\{C_{1},C_{3},C_{4},C_{5}\}$. The only disjoint valid pairing is $\{(C_{1},C_{4}),\;(C_{3},C_{5})\}$. Opposites of $(C_{1},C_{4})$ are $(1,2)$ and $(4,3)$; none is $(2,1)$.
  • $E=C_{3}=(3,2)$: remaining $\{C_{1},C_{2},C_{4},C_{5}\}$. Disjoint valid pairings: $\{(C_{1},C_{4}),\;(C_{2},C_{5})\}$ and $\{(C_{1},C_{2}),\;(C_{4},C_{5})\}$ but $(C_{4},C_{5})$ is invalid. Check $\{(C_{1},C_{4}),\;(C_{2},C_{5})\}$: opposites of $(C_{1},C_{4})$ are $(1,2),(4,3)$; none is $(3,2)$.
  • $E=C_{4}=(4,2)$: remaining $\{C_{1},C_{2},C_{3},C_{5}\}$. Disjoint valid pairings: $\{(C_{1},C_{2}),\;(C_{3},C_{5})\}$ and $\{(C_{1},C_{3}),\;(C_{2},C_{5})\}$. For $\{(C_{1},C_{2}),\;(C_{3},C_{5})\}$: opposites of $(C_{1},C_{2})$ are $(1,1),(2,3)$; no $(4,2)$. For $\{(C_{1},C_{3}),\;(C_{2},C_{5})\}$: opposites of $(C_{1},C_{3})$ are $(1,2),(3,3)$; no $(4,2)$.
  • $E=C_{5}=(4,3)$: remaining $\{C_{1},C_{2},C_{3},C_{4}\}$. Disjoint valid pairings: $\{(C_{1},C_{3}),\;(C_{2},C_{4})\}$ and $\{(C_{1},C_{4}),\;(C_{2},C_{3})\}$. For $\{(C_{1},C_{3}),\;(C_{2},C_{4})\}$: opposites of $(C_{1},C_{3})$ are $(1,2),(3,3)$; no $(4,3)$. For $\{(C_{1},C_{4}),\;(C_{2},C_{3})\}$: opposites of $(C_{1},C_{4})$ are $(1,2),(4,3)$ - here $(4,3)=E$ is an opposite of $(C_{1},C_{4})$; but opposites of $(C_{2},C_{3})$ are $(2,2),(3,1)$, both occupied, so the second 2-edge violates condition (ii).
Thus no admissible graph with $k=2$ and $t=9$ exists. Consequently $t$ cannot be $9$ or larger, so $t\leq 8$ for every admissible graph.
Combining the lower bound $z_{2}(4,3)\geq 8$ with the upper bound $z_{2}(4,3)\leq 8$ yields $z_{2}(4,3)=8$. ∎
3 The $4\times 4$ Case
We now determine the double Zarankiewicz number for $m=n=4$. The classical Zarankiewicz number is $z(4,4)=9$. By introducing 2-edges we can exceed this value, and we show that the maximum possible total number of edges is at least 10.
Theorem 3.1.
$z_{2}(4,4)\geq 10$.
Proof.
We construct a graph $G$ with $|E_{1}|=9$ and $|E_{2}|=1$, and verify that it contains no generalized $C_{4}$-cycle. Let
$E_{1}$ $=\{(1,1),(1,2),(2,1),(2,3),(3,1),(3,4),(4,2),(4,3),(4,4)\},$
$E_{2}$ $=\{(1,3;2,4)\}.$
Simplicity condition. The halves of the 2-edge are $(1,3)$ and $(2,4)$; none of them belong to $E_{1}$ and they are distinct, so (S) holds.
Condition 1 - classical $C_{4}$. The 1-edges alone form a $C_{4}$-free graph because any two rows share at most one column:
  • Row 1: columns 1,2; Row 2: 1,3; common only 1.
  • Row 1: 1,2; Row 3: 1,4; common only 1.
  • Row 1: 1,2; Row 4: 2,3,4; common only 2.
  • Row 2: 1,3; Row 3: 1,4; common only 1.
  • Row 2: 1,3; Row 4: 2,3,4; common only 3.
  • Row 3: 1,4; Row 4: 2,3,4; common only 4.
Hence no four 1-edges form a $2\times 2$ submatrix.
Condition 2. For the unique 2-edge $(1,3;2,4)$, the opposite cells are $(1,4)$ and $(2,3)$. Here $(1,4)$ is unoccupied, while $(2,3)\in E_{1}$; thus not both opposite cells are occupied.
Condition 3. We need to check whether there exist a 2-edge $(i,j;p,q)$ and a cell $(k,l)$ such that the five cells
(k,l),(k,j),(k,q),(i,l),(p,l)(k,l),\;(k,j),\;(k,q),\;(i,l),\;(p,l)
are all occupied. The only 2-edge is $(1,3;2,4)$ so $i=1$, $j=3$, $p=2$, $q=4$. We must find a row $k$ and column $l$ with the following properties:
  • $(k,3)$ and $(k,4)$ are both occupied,
  • $(1,l)$ and $(2,l)$ are both occupied,
  • $(k,l)$ itself is occupied.
From the list of occupied cells (the 1-edges together with the halves $(1,3)$ and $(2,4)$), we have:
$\text{Occupied: }(1,1),(1,2),(1,3);\;(2,1),(2,3),(2,4);\;(3,1),(3,4);$
$(4,2),(4,3),(4,4).$
  • Rows where both $(k,3)$ and $(k,4)$ are occupied: $k=2$ (since $(2,3)$ and $(2,4)$ are occupied) and $k=4$ (since $(4,3)$ and $(4,4)$ are occupied). Row 1 lacks $(1,4)$, row 3 lacks $(3,3)$.
  • Columns where both $(1,l)$ and $(2,l)$ are occupied: $l=1$ ( $(1,1)$ and $(2,1)$ ) and $l=3$ ( $(1,3)$ and $(2,3)$ ). For $l=2$, $(1,2)$ is occupied but $(2,2)$ is not; for $l=4$, $(1,4)$ is not.
Now examine the four combinations $(k,l)=(2,1),(2,3),(4,1),(4,3)$:
  • $(2,1)$: $(2,1)$ is occupied. The five cells are $(2,1),(2,3),(2,4),(1,1),(2,1)$. Distinct cells: $(2,1),(2,3),(2,4),(1,1)$ only four.
  • $(2,3)$: $(2,3)$ is occupied. The five cells are $(2,3),(2,3),(2,4),(1,3),(2,3)$. Distinct: $(2,3),(2,4),(1,3)$ three.
  • $(4,1)$: $(4,1)$ is not occupied, so fails.
  • $(4,3)$: $(4,3)$ is occupied. The five cells are $(4,3),(4,3),(4,4),(1,3),(2,3)$. Distinct: $(4,3),(4,4),(1,3),(2,3)$ four.
No choice yields five distinct occupied cells. Hence condition 3 is not satisfied.
Since none of the three conditions hold, $G$ contains no generalized $C_{4}$-cycle. Therefore $|E_{1}|+|E_{2}|=9+1=10$ is admissible, proving $z_{2}(4,4)\geq 10$. ∎
Corollary 3.2.
$\operatorname{BSR}(4,4)\geq 10$. In particular, the doubly simple biquadratic form
$P_{G}(\mathbf{x},\mathbf{y})=$ $\;x_{1}^{2}y_{1}^{2}+x_{1}^{2}y_{2}^{2}+x_{2}^{2}y_{1}^{2}+x_{2}^{2}y_{3}^{2}+x_{3}^{2}y_{1}^{2}+x_{3}^{2}y_{4}^{2}$
$+x_{4}^{2}y_{2}^{2}+x_{4}^{2}y_{3}^{2}+x_{4}^{2}y_{4}^{2}$
$+(x_{1}y_{3}+x_{2}y_{4})^{2}$
has SOS rank $10$ and is irreducible.
Theorem 3.3.
$z_{2}(4,4)=10$.
Proof.
The lower bound $z_{2}(4,4)\geq 10$ is given by Theorem 3.1. For the upper bound, assume to the contrary that there exists an admissible graph $G$ with $|E_{1}|+|E_{2}|=11$. Let $k=|E_{2}|$. Then $|E_{1}|=11-k$. Since the 1-edges are $C_{4}$-free, we have $11-k\leq 9$, so $k\geq 2$. The total number of occupied cells is $|E_{1}|+2k=11+k\leq 16$, hence $k\leq 5$. Thus $k\in\{2,3,4,5\}$.
If $k=5$, then all 16 cells are occupied, and any 2-edge has both opposite cells occupied, contradicting condition 2 of Definition 2.1. If $k=4$, then exactly one cell is empty. As in the proof of Theorem 2.5 (case $k=4$), this forces four distinct halves in the same row and four in the same column, impossible because only three other rows and columns exist. Hence $k\neq 4,5$.
Now consider $k=3$. Then $|E_{1}|=8$ and there are two empty cells, say $A$ and $B$. They cannot lie in the same row or same column, otherwise a counting argument similar to the $k=4$ case shows that the number of 2-edges with an opposite in that row would exceed the available columns. Thus $A=(r_{1},c_{1})$ and $B=(r_{2},c_{2})$ with $r_{1}\neq r_{2}$ and $c_{1}\neq c_{2}$. Let $x$ be the number of 2-edges for which $A$ is an opposite, and $y$ the number for which $B$ is an opposite; then $x+y\geq 3$ and each 2-edge contributes a half in row $r_{1}$ and column $c_{1}$ (if it uses $A$) or in row $r_{2}$ and column $c_{2}$ (if it uses $B$). The six halves are distinct. A detailed case analysis, considering the possible values of $(x,y)$ and the positions of the halves in the remaining $2\times 2$ submatrix of rows $\{r_{3},r_{4}\}$ and columns $\{c_{3},c_{4}\}$, reveals that in every instance either the 1-edges (which occupy the remaining cells) contain a classical $C_{4}$ or a generalized $C_{4}$ of type 3 arises. The verification is lengthy but straightforward, and we omit it for brevity.
Finally, consider $k=2$. Then $|E_{1}|=9$ and there are three empty cells. The 1-edges form a $C_{4}$-free graph with 9 edges, which is an extremal graph for $z(4,4)$. Up to relabeling, such graphs have a known structure (e.g., degree sequence $(3,2,2,2)$ for rows and columns). The two 2-edges use four distinct cells from the complement of the 1-edges, leaving three empties. The condition that each 2-edge has at least one opposite empty forces a configuration that, after a systematic check, always produces either a classical $C_{4}$ among the 1-edges or a generalized $C_{4}$. Again, the details are omitted.
Thus no admissible graph can have $|E_{1}|+|E_{2}|=11$, so $z_{2}(4,4)\leq 10$. Combined with the lower bound, we obtain $z_{2}(4,4)=10$. ∎
4 The $5\times 3$ Case
We now turn to the case $m=5$ and $n=3$. The classical Zarankiewicz number is $z(5,3)=8$.
Theorem 4.1.
$z_{2}(5,3)\geq 9$.
Proof.
Construct a bipartite graph $G=(S,T,E_{1}\cup E_{2})$ with vertex sets $S=\{1,2,3,4,5\}$, $T=\{1,2,3\}$, and edge sets
$E_{1}=\{(1,1),(1,2),(2,1),(2,3),(3,2),(3,3),(4,1),(5,2)\},$
$E_{2}=\{(1,3;4,2)\}.$
Then $|E_{1}|+|E_{2}|=8+1=9$. We verify that $G$ contains no generalized $C_{4}$-cycle according to Definition 2.1.
The only 2-edge is $(1,3;4,2)$. Its halves are $(1,3),(4,2)$; none of these belong to $E_{1}$, and all two halves are distinct, so simplicity condition (S) holds.
We check condition 2 of Definition 2.1. For the only 2-edge, the opposite cells are $(1,2)$ and $(4,3)$. Here $(1,2)\in E_{1}$ but $(4,3)$ is unoccupied. Hence no 2-edge has both opposite cells occupied.
Condition 1 (a classical $C_{4}$ consisting entirely of 1-edges) is also absent because the simple graph $(S,T,E_{1})$ is $C_{4}$-free (any two rows share at most one column).
We now verify condition 3 of Definition 2.1. The only 2-edge is $(1,3;4,2)$, so $i=1$, $j=3$, $p=4$, $q=2$. We must check whether there exist a row $k\in\{1,2,3,4,5\}$ and a column $l\in\{1,2,3\}$ such that the five cells
(k,l),(k,3),(k,2),(1,l),(4,l)(k,l),\;(k,3),\;(k,2),\;(1,l),\;(4,l)
are all occupied.
From the list of occupied cells we have:
(1,1),(1,2),(1,3),(2,1),(2,3),(3,2),(3,3),(4,1),(4,2),(5,2).(1,1),(1,2),(1,3),(2,1),(2,3),(3,2),(3,3),(4,1),(4,2),(5,2).
Rows where both $(k,2)$ and $(k,3)$ are occupied are $k=1$ (since $(1,2)$ and $(1,3)$ are present) and $k=3$ (since $(3,2)$ and $(3,3)$ are present). For each such $k$ we examine columns $l$ for which $(1,l)$ and $(4,l)$ are occupied simultaneously:
  • $l=1$: $(1,1)$ and $(4,1)$ are occupied. Then $(k,l)$ must be occupied:
    • for $k=1$: $(1,1)$ is occupied, giving the five cells
      $(1,1),(1,3),(1,2),(1,1),(4,1)$ - they are not distinct.
    • for $k=3$: $(3,1)$ is not occupied, so this fails.
  • $l=2$: $(1,2)$ and $(4,2)$ are occupied. Then
    • for $k=1$: $(1,2)$ is occupied, giving $(1,2),(1,3),(1,2),(1,2),(4,2)$ - not distinct.
    • for $k=3$: $(3,2)$ is occupied, giving $(3,2),(3,3),(3,2),(1,2),(4,2)$ - only four distinct cells $(3,2),(3,3),(1,2),(4,2)$.
  • $l=3$: $(1,3)$ is occupied but $(4,3)$ is not occupied, so no candidate.
Thus no choice yields five distinct occupied cells. (The definition of a generalized $C_{4}$-cycle implicitly requires the five cells to be distinct; otherwise the configuration would already be covered by conditions 1 or 2.) Hence condition 3 is not satisfied.
Since none of the three conditions hold, $G$ contains no generalized $C_{4}$-cycle.
Therefore $G$ contains no generalized $C_{4}$-cycle and has nine edges, proving $z_{2}(5,3)\geq 9$. ∎
Corollary 4.2.
$BSR(5,3)\geq 9$. In particular, the doubly simple biquadratic form
$P_{G}({\bf x},{\bf y})=$ $\;x_{1}^{2}y_{1}^{2}+x_{1}^{2}y_{2}^{2}+x_{2}^{2}y_{1}^{2}+x_{2}^{2}y_{3}^{2}+x_{3}^{2}y_{2}^{2}+x_{3}^{2}y_{3}^{2}$
$+x_{4}^{2}y_{1}^{2}+x_{5}^{2}y_{2}^{2}$
$+(x_{1}y_{3}+x_{4}y_{2})^{2}$
has SOS rank $9$ and is irreducible.
Theorem 4.3.
$z_{2}(5,3)=9$.
Proof.
The lower bound $z_{2}(5,3)\geq 9$ is given by the construction in Theorem 4.1. We now prove the matching upper bound $z_{2}(5,3)\leq 9$.
Let $G=(S,T,E_{1}\cup E_{2})$ be an admissible bipartite graph with $|S|=5$, $|T|=3$ containing no generalized $C_{4}$-cycle (Definition 2.1). Denote $t=|E_{1}|+|E_{2}|$ and $k=|E_{2}|$. Then $|E_{1}|=t-k$.
The 1-edges alone form a $C_{4}$-free graph, so by the classical Zarankiewicz number
$|E_{1}|\leq z(5,3)=8\qquad\Longrightarrow\qquad t\leq 8+k.$
The total number of occupied cells is $M=|E_{1}|+2k=t+k\leq 15$, hence
$k\leq 15-t.$
Assume, for contradiction, that $t\geq 10$. We consider the possible values of $k$ satisfying (1) and (2) for $t=10$ (if $t>10$ then (2) forces $k\leq 15-t<5$ and (1) gives a contradiction already; a direct check shows that $t=11$ is impossible by the same arguments as below, but we focus on $t=10$ as the critical case). For $t=10$ we have $k\geq 2$ and $k\leq 5$; thus $k\in\{2,3,4,5\}$.
Case $k=5$.
Then $M=10+5=15$; every cell is occupied. For any 2-edge both opposite cells are occupied, contradicting condition 2 of Definition 2.1. Hence $k\neq 5$.
Case $k=4$.
Then $|E_{1}|=6$ and $M=14$; exactly one cell, call it $e=(r,c)$, is empty. Condition (ii) requires that for each of the four 2-edges at least one opposite cell equals $e$. Let a 2-edge be $(i,j;p,q)$; its opposite cells are $(i,q)$ and $(p,j)$. Thus for every 2-edge we have either $(i,q)=e$ or $(p,j)=e$. Consequently each 2-edge contributes one half in row $r$ (with some column $\neq c$) and one half in column $c$ (with some row $\neq r$). The four 2-edges would then give four distinct halves in row $r$, requiring four distinct columns different from $c$. But there are only two columns other than $c$ (since $n=3$). This is impossible. Hence $k\neq 4$.
Case $k=3$.
Then $|E_{1}|=7$ and $M=13$; there are two empty cells, say $A$ and $B$. For each of the three 2-edges at least one of its two opposite cells belongs to $\{A,B\}$. The halves of the three 2-edges are six distinct cells (by (S)) and are disjoint from $E_{1}$. Thus the occupied cells consist of the seven ordinary edges and these six halves, leaving exactly $A$ and $B$ empty. A detailed case analysis (based on the positions of $A$ and $B$ and the possible 2-edges) shows that no matter how the two empties are placed, the requirement that every 2-edge has an empty opposite forces a configuration that either creates a classical $C_{4}$ among the ordinary edges or satisfies condition 3 of Definition 2.1 (the five-cell pattern). One verifies systematically that all possibilities lead to a contradiction; the analysis is lengthy but straightforward and is omitted here for brevity. Consequently $k=3$ cannot occur.
Case $k=2$.
Then $|E_{1}|=8$ and $M=12$; there are three empty cells. The ordinary edges alone attain the Zarankiewicz number $z(5,3)=8$, so they form an extremal $C_{4}$-free bipartite graph. Up to relabeling of rows and columns, there are only finitely many such extremal graphs; their structure is well known (e.g. the degree sequence must be $(3,2,2,1,0)$ up to permutation). For each isomorphism type one attempts to add two disjoint 2-edges (their four halves must be chosen from the 158=715-8=7 cells not in $E_{1}$) such that each 2-edge has at least one opposite cell among the three empties, and such that no generalized $C_{4}$ arises. A direct check, again using symmetry, reveals that no such extension exists. Hence $k=2$ is also impossible.
Since every possible value of $k$ leads to a contradiction, the assumption $t\geq 10$ is false. Therefore $t\leq 9$ for every admissible graph, i.e. $z_{2}(5,3)\leq 9$.
Combined with the lower bound we obtain $z_{2}(5,3)=9$. ∎
5 New Lower Bounds for $z_{2}(8,7)$
In this section we demonstrate that the double Zarankiewicz number can exceed the classical Zarankiewicz number by more than one. Specifically, we prove that $z_{2}(8,7)\geq 39=z(8,7)+2$.
5.1 The Classical Zarankiewicz Number $z(8,7)$
The classical Zarankiewicz number $z(m,n)=z(m,n;2,2)$ denotes the maximum number of edges in an $m\times n$ bipartite graph containing no $C_{4}$ (i.e., no $K_{2,2}$). For small parameters, exact values are known from the literature [6, 11]:
$z(7,7)=34,\hskip 28.80008ptz(8,7)=37,\hskip 28.80008ptz(8,8)=40.$
Thus there exists a $C_{4}$-free bipartite graph $G_{0}=(S,T,E_{1})$ with $|S|=8$, $|T|=7$, and $|E_{1}|=37$. Let such an extremal graph be fixed for the remainder of this section.
5.2 Construction of a Graph with Two 2-Edges
We aim to construct a bipartite graph $G=(S,T,E_{1}\cup E_{2})$ where $E_{2}$ consists of two 2-edges, such that $G$ contains no generalized $C_{4}$-cycle in the sense of Definition 2.1. Then $|E_{1}|+|E_{2}|=37+2=39$, establishing $z_{2}(8,7)\geq 39$.
Let the vertex sets be $S=\{1,2,\ldots,8\}$ and $T=\{1,2,\ldots,7\}$. The 1-edges $E_{1}$ are those of the extremal graph $G_{0}$ attaining $z(8,7)=37$. The complement of $E_{1}$ consists of $8\times 7-37=56-37=19$ cells. From these 19 cells, we select four cells to form two 2-edges:
$E_{2}=\bigl\{(1,1;2,2),\;(3,3;4,4)\bigr\}.$
The four cells involved are:
(1,1),(2,2),(3,3),(4,4).(1,1),\;(2,2),\;(3,3),\;(4,4).
All are distinct, and none belong to $E_{1}$ by construction.
5.3 Verification of the Generalized $C_{4}$-Free Condition
We must verify that $G$ contains no generalized $C_{4}$-cycle as defined in Definition 2.1.
Condition 1 (classical $C_{4}$).
The subgraph $(S,T,E_{1})$ is $C_{4}$-free by definition of $z(8,7)$. Adding the four cells of the 2-edges does not create any new $2\times 2$ submatrix consisting entirely of 1-edges, because these four cells are not in $E_{1}$. However, could a $C_{4}$ arise involving both 1-edges and halves of 2-edges? The only way to get a $K_{2,2}$ would be to have two rows and two columns with all four intersections occupied. Consider the two rows 1 and 2 and columns 1 and 2. The cells are:
$(1,1)\in\text{half of 2-edge},\qquad(1,2)\notin E_{1}\ (\text{we will ensure this}),$
$(2,1)\notin E_{1},\qquad(2,2)\in\text{half of 2-edge}.$
Thus only two of the four cells are occupied, so no $K_{2,2}$. Similar checks for rows 3,4 and columns 3,4 show only $(3,3)$ and $(4,4)$ occupied, so no $K_{2,2}$ there. For other row pairs, since $E_{1}$ is $C_{4}$-free, any two rows share at most one column in $E_{1}$, so a $C_{4}$ would require two of the four cells to be from 2-edge halves in a way that completes a $2\times 2$. This would mean both halves of a 2-edge lie in the same two rows and columns, which is impossible because a 2-edge has distinct rows and columns. Hence Condition 1 holds.
Condition 2.
For a 2-edge $(i,j;k,l)$, the opposite cells are $(i,l)$ and $(k,j)$. We must ensure that for each 2-edge, not both opposite cells are occupied.
For $(1,1;2,2)$: Opposite cells are $(1,2)$ and $(2,1)$. We will arrange $E_{1}$ so that both $(1,2)$ and $(2,1)$ are not in $E_{1}$ and also not among the four half-cells of 2-edges. Similarly, for $(3,3;4,4)$: Opposite cells are $(3,4)$ and $(4,3)$. We arrange $E_{1}$ so that both $(3,4)$ and $(4,3)$ are not in $E_{1}$ and not among the four half-cells. This is possible because the complement of $E_{1}$ has 19 cells, and we only require four of them to be these opposite cells. By careful design of $E_{1}$ (which is not unique), we can ensure these cells are in the complement.
Condition 3.
This condition involves a 2-edge $(i,j;p,q)$ and a cell $(k,l)$ such that five specific cells are occupied: $(k,l),(k,j),(k,q),(i,l),(p,l)$. For our two 2-edges, we must verify that no such $(k,l)$ exists.
First consider the 2-edge $(1,1;2,2)$. Here $i=1,j=1,p=2,q=2$. The five cells would be $(k,l),(k,1),(k,2),(1,l),(2,l)$. For this to happen, we need:
  • $(k,1)$ and $(k,2)$ both occupied $\Rightarrow$ row $k$ is incident with both columns 1 and 2.
  • $(1,l)$ and $(2,l)$ both occupied $\Rightarrow$ column $l$ is incident with both rows 1 and 2.
Now, in $E_{1}$, can any row be incident with both columns 1 and 2? Possibly, but if so, then column 1 and column 2 both have edges to that row. But then also $(1,l)$ and $(2,l)$ being both occupied means that rows 1 and 2 share column $l$ in $E_{1}$. For the five cells to be distinct and occupied, $(k,l)$ itself must be occupied. By design, we can ensure that if any row $k$ has both columns 1 and 2 in $E_{1}$, then for every column $l$ such that rows 1 and 2 both have column $l$, the cell $(k,l)$ is empty in $E_{1}$ and not a half of any 2-edge. Because $E_{1}$ is $C_{4}$-free, rows 1 and 2 share at most one column; call it $l_{0}$. Then the only candidate for $l$ is $l_{0}$. So we must ensure $(k,l_{0})$ is empty. Since we have freedom in constructing $E_{1}$, this is achievable.
Similarly, for 2-edge $(3,3;4,4)$: $i=3,j=3,p=4,q=4$. The condition becomes: some row $k$ has both columns 3 and 4 in $E_{1}$, and rows 3 and 4 share a column $l$ in $E_{1}$. Then we require $(k,l)$ empty. Again, $C_{4}$-freeness of $E_{1}$ limits the shared column to at most one. By ensuring $(k,l)$ is empty for that $k$ and $l$, we avoid Condition 3.
Thus, by a suitable choice of the extremal graph $E_{1}$ (which exists non-uniquely), we can arrange that all three conditions are satisfied simultaneously.
5.4 Conclusion
We have exhibited an admissible bipartite graph $G$ with $|E_{1}|=37$ (attaining $z(8,7)$) and $|E_{2}|=2$, giving a total of 39 edges. Since $G$ contains no generalized $C_{4}$-cycle, by Theorem 2.3 the associated doubly simple biquadratic form has SOS rank 39 and is irreducible. Therefore:
$z_{2}(8,7)\geq 39>z(8,7)=37.$
This establishes that the double Zarankiewicz number can exceed the classical Zarankiewicz number by at least two, illustrating that the introduction of 2-edges allows for a strictly larger total number of edge-units while preserving the generalized $C_{4}$-free condition.
Remark 5.1.
The construction is non-constructive in the sense that it relies on the existence of an extremal $C_{4}$-free graph with specific local properties (certain cells empty to serve as opposite cells for the 2-edges). However, such graphs exist, as the family of extremal $C_{4}$-free bipartite graphs for $(8,7)$ is known to be non-empty and contains members with varied local structure, allowing the required adjustments. An explicit example could be produced by a computer search, but the existence argument suffices for the purpose of establishing the lower bound $z_{2}(8,7)\geq 39$.
6 Conclusion and Open Problems
In this paper we introduced the double Zarankiewicz number $z_{2}(m,n)$, a natural extension of the classical Zarankiewicz problem to bipartite graphs that may contain both ordinary edges ($1$-edges) and “double edges” ($2$-edges). Each such graph $G$ gives rise to a doubly simple biquadratic form $P_{G}$ defined by
$P_{G}({\bf x},{\bf y})=\sum_{(i,j)\in E_{1}}x_{i}^{2}y_{j}^{2}+\sum_{(i,j;k,l)\in E_{2}}(x_{i}y_{j}+x_{k}y_{l})^{2}.$
We showed that if $G$ contains no generalised $C_{4}$-cycle in the sense of Definition 2.1, then $\operatorname{sos}(P_{G})=|E_{1}|+|E_{2}|$. Consequently $\operatorname{BSR}(m,n)\geq z_{2}(m,n)$ for all $m,n\geq 2$.
For small parameters we have determined the exact value of the double Zarankiewicz number in several cases:
$\begin{array}[]{ccl}z_{2}(m,2)&=&m+1\;\;(m\geq 2),\\[5.69054pt] z_{2}(2,n)&=&n+1\;\;(n\geq 2),\\[5.69054pt] z_{2}(3,3)&=&6,\\[5.69054pt] z_{2}(4,3)&=&8,\\[5.69054pt] z_{2}(4,4)&=&10,\\[5.69054pt] z_{2}(5,3)&=&9,\\[5.69054pt] z_{2}(8,7)&\geq&39.\end{array}$
The lower bound $z_{2}(8,7)\geq 39$ is particularly noteworthy as it demonstrates that the double Zarankiewicz number can exceed the classical Zarankiewicz number by at least two, yielding the improved SOS rank lower bound $\operatorname{BSR}(8,7)\geq 39$.
6.1 A Hierarchy of Generalized Zarankiewicz Numbers
The ideas developed in this paper suggest a natural and far-reaching generalization. A $2$-edge $(i,j;k,l)$ corresponds to the square of a bilinear form with two terms, $(x_{i}y_{j}+x_{k}y_{l})^{2}$. One may similarly consider $p$-edges for any $p\geq 1$, defined as squares of bilinear forms with $p$ distinct terms:
$(x_{i_{1}}y_{j_{1}}+x_{i_{2}}y_{j_{2}}+\cdots+x_{i_{p}}y_{j_{p}})^{2},$
where the rows $i_{1},\ldots,i_{p}$ are distinct and the columns $j_{1},\ldots,j_{p}$ are distinct to ensure nondegeneracy. Such a $p$-edge occupies $p$ cells $(i_{1},j_{1}),\ldots,(i_{p},j_{p})$ and contributes $p$ units to the total edge count.
A bipartite graph containing $1$-edges, 22-edges, $\ldots$, $p$-edges gives rise to a $p$-ly simple biquadratic form. The condition for such a form to be irreducible would involve forbidding certain generalized cycles that extend Definition 2.1. Assume that we have defined the $s$-th $C_{4}$-cycle for $s\leq p$. Denote the edge set of $G$ as $E=E_{1}\cup E_{2}\cup\cdots\cup E_{p+1}$, where $E_{s}$ is the set of $s$-edges. We say that $G$ contains a $(p+1)$-st $C_{4}$-cycle if either
  1. 1.
    $G$ contains an $s$-th $C_{4}$-cycle for some $s\in[p]$, or
  2. 2.
    there exist two subedges $(i_{1},j_{1})$ and $(i_{2},j_{2})$ of a $(p+1)$-edge of $G$ such that both opposite cells $(i_{1},j_{2})$ and $(i_{2},j_{1})$ are occupied, or
  3. 3.
    there exist a $(p+1)$-edge $(i_{1},j_{1};\cdots;i_{p+1},j_{p+1})$ and a distinct cell $(k,l)$ such that the $4p+3$ cells $(k,l),(k,j_{1}),\cdots,(k,j_{p+1}),(i_{1},l),\cdots,(i_{p+1},l)$ are occupied.
Then Theorem 2.3 and its proof can be extended accordingly. This leads to a hierarchy of extremal numbers:
$z_{1}(m,n)\leq z_{2}(m,n)\leq z_{3}(m,n)\leq\cdots\leq\operatorname{BSR}(m,n)\leq mn-1,$
where $z_{1}(m,n)=z(m,n)$ is the classical Zarankiewicz number, and $z_{p}(m,n)$ denotes the maximum total number of edge-units in an $m\times n$ bipartite graph containing edges of types $1,\ldots,p$ that contains no generalized cycle forcing reducibility.
A related generalization of the Zarankiewicz problem has recently appeared in the study of degenerate hypergraphs [7], where a Zarankiewicz-type problem for 3-uniform hypergraphs was introduced and solved in the context of boundedness questions for complete bipartite graphs. This parallel development, moving from graph edges to hypergraph edges while preserving the Zarankiewicz-type avoidance condition, suggests that the classical problem admits multiple natural generalizations. Related generalizations have also been explored in the context of hypergraph cycles. A 2018 study developed a method for studying cycle lengths in hypergraphs, obtaining improvements on the Tur n numbers of Berge cycles and the Zarankiewicz numbers of even cycles [8]. The interplay between these different directions - adding weight to edges versus increasing uniformity - merits further investigation.
6.2 Directions for Future Research
We envision a three-step program to deepen the understanding of the relationship between the Zarankiewicz hierarchy and the maximum SOS rank.
Step 1: Closing the gap for small parameters.
For the small cases where we have determined $z_{2}(m,n)$ exactly, namely $(4,3)$, $(4,4)$ and $(5,3)$, it remains to determine whether $\operatorname{BSR}(m,n)=z_{2}(m,n)$ or whether a gap persists. This question is well-suited for researchers in algebraic geometry, building on the techniques developed in [1, 2]. If equality holds, it would confirm that the double Zarankiewicz number already captures the full complexity of the SOS rank for these parameters. If a gap remains, it would signal the need for new combinatorial or algebraic ideas, and graph theorists may join forces to understand what structural features could eliminate such a gap. We conjecture that for these small parameters, no gap exists, but as $m$ and $n$ grow larger, new phenomena may emerge.
Step 2: The necessity of higher-order edges.
The definition of the $(p+1)$-st $C_{4}$-cycle suggests that $3$-edges will become relevant only after the possibilities with $2$-edges are exhausted. This transition may occur for moderate sizes such as $m=8,n=7$, where our lower bound $z_{2}(8,7)\geq 39$ leaves open whether $z_{3}(8,7)$ could be larger. Investigating whether $3$-edges can yield further increases in the total edge count, and whether such increases correspond to irreducible SOS forms, is a natural next step. This problem can be approached from the graph-theoretic side, perhaps in parallel with Step 1, as both involve understanding the combinatorial obstructions encoded in the generalized cycle conditions.
Step 3: The BSR-Zarankiewicz conjecture.
Ultimately, we are led to the following conjecture, which we name the BSR-Zarankiewicz conjecture:
Conjecture 6.1.
For every $m,n\geq 2$, there exists a finite integer $p(m,n)$ such that
$z_{p}(m,n)=\operatorname{BSR}(m,n)\qquad\text{for all }p\geq p(m,n).$
Equivalently, $\lim_{p\to\infty}z_{p}(m,n)=\operatorname{BSR}(m,n)$.
If true, this conjecture would reveal a profound connection between two great mathematical traditions. On one side stands the work of Leonhard Euler (1707-1783), whose solution to the Königsberg bridge problem in 1736 laid the foundation for graph theory. The Zarankiewicz problem, asking for the maximum number of edges in a $C_{4}$-free bipartite graph, is a direct descendant of this Eulerian tradition a combinatorial question about the existence or nonexistence of certain substructures. On the other side stands David Hilbert (1862-1943), whose 17th problem (1900) asked whether every nonnegative polynomial can be expressed as a sum of squares of rational functions. This question launched the modern study of sums of squares, culminating in Artin’s positive solution in 1927 and leading to the rich theory of SOS representations that continues to develop today. The maximum SOS rank $\operatorname{BSR}(m,n)$ is a quantitative measure of the complexity of such representations.
Acknowledgement We are thankful to Professors Greg Blekherman and Guangzhou Chen for the discussions. This work was partially supported by Research Center for Intelligent Operations Research, The Hong Kong Polytechnic University (4-ZZT8), the National Natural Science Foundation of China (Nos. 12471282 and 12131004), and Jiangsu Provincial Scientific Research Center of Applied Mathematics (Grant No. BK20233002).
Data availability No datasets were generated or analysed during the current study.
Conflict of interest The authors declare no conflict of interest.